### v1

# gamecenter,boxed

Abstract.
We study the problem of privacy-preserving $k$-means clustering in the horizontally federated setting. Existing federated approaches using secure computation, suffer from substantial overheads and do not offer output privacy. At the same time, differentially private (DP) $k$-means algorithms assume a trusted central curator and do not extend to federated settings. Naively combining the secure and DP solutions results in a protocol with impractical overhead. Instead, our work provides enhancements to both the DP and secure computation components, resulting in a design that is faster, more private, and more accurate than previous work. By utilizing the computational DP model, we design a lightweight, secure aggregation-based approach that achieves four orders of magnitude speed-up over state-of-the-art related work. Furthermore, we not only maintain the utility of the state-of-the-art in the central model of DP, but we improve the utility further by taking advantage of constrained clustering techniques.
Differential Privacy, Federated Learning, $k$-Means Clustering, Secure Computation
ccs: Computing methodologies Cluster analysisccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Data anonymization and sanitization
1. Introduction
Unsupervised learning allows data analysts to extract meaningful patterns from raw data that may be difficult to label. The canonical example of unsupervised learning is the Euclidean $k$-means clustering problem, where data is grouped into clusters with similar features. Clustering has a plethora of important applications in recommendation systems, fraud detection, and healthcare analytics. In such applications, the dataset often contains sensitive data, which necessitates the use of privacy-preserving techniques. Combining data from multiple parties in the federated scenario yields more robust insights about the global population at the cost of further exasperating the privacy risks.
To counteract the privacy risks, a number of works focus on solving the federated $k$-means problem using secure computation (Vaidya and Clifton, 2003; Jha et al., 2005; Jagannathan and Wright, 2005; Jagannathan et al., 2010; Gheid and Challal, 2016; Bunn and Ostrovsky, 2007; Jäschke and Armknecht, 2018; Mohassel et al., 2020; Patel et al., 2012; Yuan and Tian, 2019; Rao et al., 2015; Xiaoyan et al., 2015; Silva and Bellala, 2017; Jiang et al., 2020). Secure computation enables analysts to solve the $k$-means problem while keeping the sensitive data encrypted. However, all of these works still output the exact result of the clustering. This gives a false sense of privacy; although the input and intermediate computations all remain private, no effort is made to protect the privacy of the output. It is well known in the literature that publishing exact statistics can leak significant information about the input data and allow attacks such as dataset reconstruction (Dwork et al., 2017). This level of privacy does not justify the significant runtime overhead incurred by solely using secure computation.
A separate line of research considers using differential privacy (DP) to protect the privacy of the output of $k$-means (Blum et al., 2005; Dwork, 2011; Su et al., 2016; Mohan et al., 2012; Zhang et al., 2013; Balcan et al., 2017; Lu and Shen, [n. d.]; Park et al., 2017; Ni et al., 2021). Differential privacy guarantees that the clustering’s output will be approximately the same regardless of any individual user’s participation. However, existing methods do not consider the federated setting and assume the existence of a trusted central party who runs the $k$-means algorithm and sanitizes the results with noise before publishing.
To summarize, no prior work offers a solution to protect both the input and output privacy of the federated $k$-means problem. A straw man solution to our problem would be to combine the current state-of-the-art solutions in secure federated $k$-means and DP $k$-means. However, this solution would be unacceptable as the runtime of current federated $k$-means approaches using secure computation is prohibitively large. The state-of-the-art approach of Mohassel et al. (Mohassel et al., 2020) incurs runtime overheads on the order of tens of minutes. Naively adding differentially private noise to this computation will further increase this runtime.
Our work focuses on designing an efficient, DP $k$-means in the federated setting that protects the privacy of the input, output, and intermediate computations. Following related work in federated $k$-means (Mohassel et al., 2020; Jiang et al., 2020), we use Lloyd’s algorithm as the foundation of our protocol. However, we design a new variant of DP-Lloyd (Blum et al., 2005; Su et al., 2016) that noises the centroids directly. Our algorithm improves the clustering utility over the state-of-the-art DP algorithm of Su et al. (Su et al., 2016) by adding constraints to the $k$-means objective (Bennett et al., 2000). Specifically, by adding minimum and maximum constraints to the cluster size, we achieve a non-trivial bound on the sensitivity of the centroids. This enables our algorithm to aggregate at the centroid level and apply less DP noise than prior work that aggregates the sums and counts separately (Su et al., 2016).
To tailor our DP-Lloyd algorithm to secure computation, we design a protocol that publishes the centroids in each step. This allows operations that would be expensive in secure computation (such as assignment and division) to be conducted in plain text. However, unlike prior work (Gheid and Challal, 2016; Jiang et al., 2020), we ensure that the intermediate centroids are protected by DP guarantees. We prove that the protocol is secure in the computational DP definition (Mironov et al., 2009), which allows DP-bounded leakage during the computation. Typically, utilizing the computational DP model to leak intermediate steps implies sacrificing privacy or utility for an efficient runtime. However, in our case, we specifically choose to leak intermediate values that are typically accounted for in the central DP proof of DP-Lloyd (Blum et al., 2005; Su et al., 2016). In other words, we get this speed-up with no additional cost to the privacy parameter $\epsilon$. Furthermore, rather than decrease the utility, we improve the utility of the clustering over state-of-the-art approaches (Su et al., 2016).
To perform the aggregation over multiple parties, we design a modified secure aggregation protocol that keeps the aggregator oblivious to the client’s inputs and the centroids that are output at each iteration of the protocol. This allows us to add DP noise equivalent to the central model of DP. The combination of our improved DP algorithm, leaking the DP centroids, and our lightweight secure aggregation protocol allows us to reduce the computation time by up to four orders of magnitude compared to related work (Mohassel et al., 2020) (from minutes to a fraction of a second per iteration).
In summary, our contributions include:
  • We design the first DP protocol for horizontally federated private $k$-means.
  • We improve the clustering utility over state-of-the-art DP $k$-means solutions by using cluster-size constraints and centroid-level aggregation.
  • We design an efficient protocol using DP and a modified secure aggregation protocol to implement our protocol in the local model of trust.
  • We prove our protocol is secure and preserves the end-to-end privacy in the computational model of DP and reduces the runtime by four orders of magnitude over the state-of-the-art secure federated approaches.
The remainder of the paper is organized as follows. We begin with some problem-specific background information in Section 2 and formally define the problem in Section 3. In Section 4, we summarize the relevant literature. We then present our complete protocol for federated DP $k$-means (FastLloyd) and prove its privacy and utility in Section 5. In Section 6, we give an in-depth evaluation of our protocol in terms of utility, runtime, and communication size over various real-world and synthetic datasets. Finally, Section 7 describes extensions to our protocol, including a distance-constrained method and alternatives to the aggregation server.
2. Background
2.1. Notation
Throughout this paper, we will denote objects that we intend to further slice/index with boldface notation (e.g. $\boldsymbol{\mu}$, while keeping atomic objects (whether scalars or vectors) as normal face (e.g. $\epsilon$). A summary of the notation used in this paper is provided in Table 1 and we will define the notation as it is used.
Symbol Description
$\boldsymbol{p}$ Set of $M$ clients: $\{p_{1},p_{2},...,p_{M}\}$
S Service provider
$D_{i}$ Local dataset of client $p_{i}$: $\{x_{1},x_{2},...,x_{N_{i}}\}$
$N_{i}$ Size of the dataset of client $p_{i}$
$d$ Dimension of the data points
$k$ Number of clusters
$\mathbf{O}$ Set of $k$ clusters: $\{O_{1},O_{2},...,O_{k}\}$
$\boldsymbol{\mu}$ Centroids of clusters: $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
$\boldsymbol{\gamma}$ Differentially-private noise
$\eta$ Constraint parameter
$\epsilon$ Privacy budget
Table 1. Notation used in the paper.
2.2. $k$-means Lloyd’s Algorithm
The $k$-means problem is a discrete optimization problem that aims to partition a set of $N$ $d$-dimensional observations into $k$ clusters, each represented by its mean or centroid. Given a dataset of observations $D=\{x_{1},x_{2},...,x_{N}\}$, where each observation is a $d$-dimensional real vector, the $k$-means problem is to find an assignment of data points to clusters, and a set of cluster centers, that minimizes the Within-cluster Sum of Squares (WCSS) objective:
(1) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
where $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$, $O_{\subseteq}D$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the centroids of the clusters, defined to be the (arithmetic) mean of points in $O_{j}$.
Lloyd’s Algorithm
The $k$-means problem is NP-hard in Euclidean space (Aloise et al., 2009). However, a simple heuristic known as Lloyd’s algorithm is commonly applied for practical applications (Lloyd, 1982). We detail the basic procedure of Lloyd’s algorithm:
  1. (1)
    Initialization step: Randomly sample $k$ centroids $\{\mu_{1},\mu_{2},...,\mu_{k}\}$ (typically from the datapoints).
  2. (2)
    Repeat until convergence (the assignments no longer change) or a predetermined number of iterations has been reached:
    1. (a)
      Assignment step: Assign each observation $x_{l}$ to the nearest centroid (using the Euclidean distance). This creates clusters $O_{j}$ for $j=1,2,...,k$. Formally, the assignment is:
      $O_{j}^{(t)}=\big{\{}x_{l}:||x_{l}-\mu_{j}^{(t-1)}||^{2}\leq||x_{l}-\mu_{c}^{(t-1)}||^{2}\ \forall c\in\{1,\dots,k\}\}$
    2. (b)
      Update step: Calculate the new centroids to be the mean of the observations in the cluster:
      $\mu_{j}^{(t)}=\frac{\sum\limits_{x_{l}\in\mathbf{O}{j}^{(t)}}x_{l}}{\left|O_{j}^{(t)}\right|}$
2.3. Constrained $k$-means
The constrained $k$-means problem is a variant of the $k$-means problem where, in addition to minimizing the WCSS (Eqn 1), a certain set of constraints $\boldsymbol{\eta}$ must be satisfied. The problem was first introduced by Bennet et. al (Bennett et al., 2000), where they propose a method to enforce minimum cluster size constraints, $\boldsymbol{\eta}^{min}$.
To solve the constrained $k$-means problem, Lloyd’s algorithm can be modified to incorporate the constraints in the assignment step of every iteration. However, the assignment step becomes nontrivial as it requires the solution to a constrained optimization problem. Bennet et. al (Bennett et al., 2000) provide an efficient solution to this constrained optimization problem by formulating it as a min-cost flow problem, which can be solved using off-the-shelf network flow solvers. An open-source implementation by Levy-Kramer (Levy-Kramer, 2018) implements the approach, using Google’s ORTools (Perron and Furnon, 2023) to solve the flow problem, and extends it to allow for maximum cluster size constraints, $\boldsymbol{\eta}^{max}$ (Eqn 2).
(2) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}\quad\text{s.t.}\quad\eta_{j}^{min}\leq|O_{j}|\leq\eta_{j}^{max}\quad\forall j$
Figure 1 shows an example reduction of an assignment instance to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
[…]
Figure 1. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
Datapoints are represented as nodes and connected to cluster nodes through directed edges. The cost of each edge $(x_{l}\rightarrow O_{j})$ is defined to be the Euclidean distance between $x_{l}$ and the corresponding centroid $\mu_{j}$. The capacity of each edge is set to $1$, and each data node has a supply of $1$, ensuring that each data point is associated with exactly one cluster. On the other hand, each centroid node has a supply of $-\eta_{min}$ (equivalently, a demand of $\eta_{min}$). As every valid flow solution must satisfy this demand, a minimum of $\eta_{min}$ data points will flow into each cluster. To impose the maximum size constraint, clone nodes are created for each centroid, and each centroid is connected with its clone through an edge with zero cost and a capacity of $\eta_{max}$. This constrains the maximum flow into each cluster to be $\eta_{max}$. Lastly, a sink node is introduced to consume the remaining flow. A valid flow solution respects the constraints on cluster sizes, and by minimizing the cost, the total distance of points to their assigned centroids is minimized. Therefore, a min-cost flow solution to this model solves the constrained $k$-means problem for both minimum and maximum cluster constraints.
2.4. Differential Privacy
Differential privacy (DP) (Dwork et al., 2006) is an increasingly popular notion for protecting individuals’ privacy while allowing the computation of aggregate statistics. Differential privacy guarantees that an algorithm’s output is approximately the same, regardless of the participation of any single user. More formally, differential privacy can be defined as follows.
Definition 2.1 (Differential Privacy).
A randomized algorithm $M:\mathcal{D}\mapsto\mathbb{R}$ is $\epsilon$-DP, if for any pair of neighbouring datasets $D,D^{\prime}\in\mathcal{D}$, and for any $T\subseteq\mathbb{R}$ we have
(3) $\Pr[M(D)\in T]\leq e^{\epsilon}\Pr[M(D^{\prime})\in T].$
The privacy parameter $\epsilon$ defines how similar the outputs must be and determines the trade-off between privacy and utility. We use the unbounded neighbouring definition where datasets are neighbours if $|D\backslash D^{\prime}\cup D^{\prime}\backslash D|=1$. That is, we allow the addition or removal of a single data point. We note that arbitrary computations can be carried out on the output of a DP mechanism without affecting privacy (the postprocessing lemma (Dwork and Roth, 2014)). Finally, DP is composed naturally with multiple runs of a mechanism. If we apply a differentially private mechanism(s) sequentially, the privacy parameters are composed through summation or more advanced methods (Dwork and Roth, 2014). If a mechanism is applied multiple times over disjoint subsets of the dataset, then the total privacy leakage is the maximum privacy parameter over each subset (parallel composition (Dwork and Roth, 2014)).
Definition 2.2 (Sensitivity).
Let $f:\mathcal{D}\mapsto\mathbb{R}^{k}$. If $\mathbb{D}$ is a distance metric between elements of $\ \mathbb{R}^{k}$ then the $\mathbb{D}$-sensitivity of $f$ is
(4) $\Delta^{(f)}_{d}=\max_{(D,D^{\prime})}\mathbb{D}(f(D),f(D^{\prime})),$
where $(D,D^{\prime})$ are pairs of neighbouring datasets.
Common examples of the distance metric are the $\ell_{1}$-norm and $\ell_{2}$-norm. With the notion of sensitivity, we can now define the Laplace mechanism, which uses sensitivity based on the $\ell_{1}$-norm, which we denote $\Delta^{(f)}_{1}$ for simplicity.
Definition 2.3 (Laplace Mechanism).
Let $f:\mathcal{D}\mapsto\mathbb{R}^{k}$. The Laplace mechanism is defined as
(5) $M(D)=f(D)+\boldsymbol{\gamma},$
where
$\boldsymbol{\gamma}\in\mathbb{R}^{k}\quad\text{s.t.}\quad\gamma_{i}\overset{\mathrm{iid}}{\sim}Lap(\Delta^{(f)}_{1}/\epsilon)$
i.e. $\boldsymbol{\gamma}$ represents $k$ i.i.d draws from the Laplace distribution with scale parameter $\Delta^{(f)}_{1}/\epsilon$.
It is well known that the Laplace mechanism satisfies $\epsilon$-DP (Dwork and Roth, 2014).
3. Problem Statement
We consider $M$ clients denoted by $\boldsymbol{p}=\{p_{1},p_{2},...,p_{M}\}$, with each party $p_{i}$ owning a private dataset $D_{i}=\{x_{1},x_{2},...,x_{N_{i}}\}$, where $x_{l}\in\mathbb{R}^{d}$, $d$ denotes the dimensionality of the dataset and $N_{i}$ is the size of the dataset of party $i$. For simplicity, we assume that datasets held by each client are of similar sizes. The objective is to compute a collaborative $k$-means clustering $(\mathbf{O},\boldsymbol{\mu})$ over $\boldsymbol{D}$ that minimizes the WCSS objective (1); where $\boldsymbol{D}=\bigcup_{i=1}^{i=M}D_{i}$ is the union of the datasets held by the clients, $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the respective centroids of the clusters.
We aim to optimize this objective while formally proving the privacy of the output and intermediate computations. Specifically, the protocol should be secure in the computational model of differential privacy introduced by Mirnov et al. (Mironov et al., 2009) and extended to the multi-party setting by Humphries et al. (Humphries et al., 2022).
Definition 3.1 (IND-CDP-MPC (Humphries et al., 2022; Mironov et al., 2009)).
A multi-party protocol $\Pi$ for computing function $f$ satisfies $\epsilon(\lambda)$- indistinguishable computationally differential privacy (IND-CDP-MPC) if for every probabilistic polynomial time (in $\lambda$) adversary $\mathcal{A}$ with input $D_{A}$, and for neighbouring datasets $D,D^{\prime}$ belonging to the honest parties (i.e. $D,D^{\prime}=\cup_{i\setminus A}D_{i}$),
$\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D))=1]$
$\leq$ $\exp(\epsilon)\cdot\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D^{\prime}))=1]+negl(\lambda).$
Likewise, for every other party’s view of neighbours $(D,D^{\prime})$.
Intuitively, after observing the output and intermediate computations, the amount of information that any party can learn about any other party’s local dataset should be bounded by differential privacy.
Finally, we make the following assumptions when designing our protocol:
  • The existence of an honest-but-curious service provider S to assist with multiparty computations. This party should be oblivious to all inputs and results, i.e., the service provider learns nothing about the client’s input or output during the execution of the protocol. We discuss alternatives to this assumption in Section 7.
  • Clients share a common secret before the start of the protocol; this can be used as a seed to a common Pseudo-Random Number Generator (PRNG) and is generally possible through public-key schemes or communication over a channel secured by TLS.
4. Related Work
4.1. Secure Exact $k$-means
We first discuss related work using secure computation to compute $k$-means that does not preserve the privacy of the output. This literature shows a trade-off between privacy and efficiency. Approaches that prioritize efficiency often compromise privacy by allowing the leakage of intermediate computations such as sums, counts, or centroids, which can be exploited to infer sensitive data (Gheid and Challal, 2016; Jiang et al., 2020). For instance, Gheid et al. (Gheid and Challal, 2016) rely on computing secure sums for aggregating sums and counts across clients. While this is very efficient in practice, it reveals aggregate intermediate values to all clients, which is not considered secure (Hegde et al., [n. d.]). Another approach by Jiang et al. (Jiang et al., 2020) employs homomorphic encryption and garbled circuits to conceal sums and counts. However, it still reveals the intermediate centroids to all clients in the clear. Many other works in this category are explained in the SoK of Hedge et al. (Hegde et al., [n. d.]).
Conversely, secure approaches that do not leak intermediate computations employ heavy cryptographic techniques, like fully homomorphic encryption, leading to significant performance degradation (Bunn and Ostrovsky, 2007; Rao et al., 2015; Jäschke and Armknecht, 2018; Kim and Chang, 2018). Bunn and Ostrovsky (Bunn and Ostrovsky, 2007) utilize arithmetic secret sharing alongside homomorphic encryption, with high-performance overheads. Homomorphic encryption has also been used by Rao et al. (Rao et al., 2015), Jäschke et al. (Jäschke and Armknecht, 2018), and Kim et al. (Kim and Chang, 2018), and consistently results in prohibitively high computation time due to the computation demands of the encryption. The work of Mohassel et al. (Mohassel et al., 2020) is a notable exception, offering a scheme that is significantly faster than previous state-of-the-art methods. By using optimized $2$-party computation primitives, batched oblivious transfers, and garbled circuits, they achieve a speedup of five orders of magnitude over Jäschke et al. (Jäschke and Armknecht, 2018). However, the runtime is still on the order of minutes; and the communication size is on the order of gigabytes. In contrast, our approach is as efficient as the protocols that leak intermediate computations, while protecting the privacy of the output and intermediate computations with DP.
4.2. Differentially Private $k$-means
DP $k$-means algorithms aim to address output privacy by introducing noise during clustering, hiding individual data contributions, and ensuring a formal privacy guarantee. Research in DP $k$-means is divided into central and local models.
Central DP
The central model operates with a trusted curator who collects data for analysis, hiding individuals’ information in the output, but not from the curator itself. Among the central approaches, DP-Lloyd introduces noise during each centroid update, by adding Laplace noise to both the sum (numerator) and the counts (denominator) when computing the arithmetic mean of points within a cluster (Blum et al., 2005). DP-Lloyd was later implemented in the PinQ framework (McSherry, 2009) with a fixed number of iterations. Dwork (Dwork, 2011) extended this framework to allow for an arbitrary number of iterations, allocating an exponentially decreasing privacy budget to each iteration, although it was noted that utility degrades beyond a certain number of iterations due to increasing noise levels. The state-of-the-art in DP-Lloyd was achieved by Su et al. (Su et al., 2016), who included an error analysis to determine the optimal number of iterations and introduced a “sphere-packing” centroid initialization method that significantly improves clustering quality.
The Sample and Aggregate Framework (SaF) partitions the dataset into multiple subsets, upon which the non-private Lloyd’s algorithm is executed independently (Nissim et al., 2007). The resulting centroids from each subset are then aggregated using a standard DP mechanism. SaF was later implemented within the GUPT system (Mohan et al., 2012). However, experiments conducted by Su et al. (Su et al., 2016) suggest that DP-Lloyd consistently outperforms SaF across a variety of synthetic and real datasets. Synopsis-based methods take a different approach by first generating a synopsis of the dataset using a differentially private algorithm, then applying the $k$-means clustering algorithm on this synopsis. Qardaji et al. (Qardaji et al., 2012) proposed a synopsis technique for 2D datasets, which was extended and optimized by Su et al. (Su et al., 2016) to higher dimensions. However, this approach only outperforms DP-Lloyd in datasets with less than three dimensions (Su et al., 2016). Another, more theoretical, line of work focuses on minimizing approximation error bounds, without implementation or empirical evaluation (Feldman et al., 2009, 2017; Balcan et al., 2017; Nissim and Stemmer, 2018; Stemmer and Kaplan, 2018; Ghazi et al., 2020). The work of Ghazi et al. (Ghazi et al., 2020) concludes this long line of work by achieving the same approximation factor as the best non-private algorithms. However, these methods suffer from superlinear running times, making them impractical for large datasets, where privacy is most important.
Local DP
In contrast to central DP, the local model operates under the premise that no trusted curator is available. This necessitates that individuals randomize their data locally before it is aggregated, which reduces utility by introducing a significantly higher level of noise (on the order of the number of clients). Most research in this area, including works by Nissim and Stemmer (Nissim and Stemmer, 2018) and Stemmer (Stemmer, 2021), focus on theoretical error bounds rather than practical applications. Finally, Li et al. (Li et al., 2023) attempt to bridge the gap between these central and local models in the vertically federated setting. Their approach relies on an untrusted curator aggregating noisy local centers and membership encodings from the clients to generate a private synopsis. However, this approach is specifically tailored for the vertical setting and cannot be applied to our configuration. Our work yields utility that is even better than state-of-the-art work in the central model (Su et al., 2016) while obtaining the trust assumptions of the local model and operating in the federated setting.
5. Distributed DP $k$-means
First, we overview our modified DP algorithm, using constrained $k$-means. Then, we describe FastLloyd, including the initialization, assignment, and update steps. Finally, we provide an error analysis and proof of privacy.
5.1. Constrained DP $k$-means
We observe that neither federated exact $k$-means nor DP $k$-means algorithms in the literature take advantage of constrained clustering approaches from the non-private literature (Bennett et al., 2000; Levy-Kramer, 2018). In this section, we describe how incorporating constrained clustering into the assignment step significantly improves utility; a contribution that is independently applicable to the central model of DP.
In the DP $k$-means literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), the typical way to compute a differentially-private mean is to perform two differentially-private queries: one for the sum and one for the count. The sum is then divided by the count as a postprocessing step. We call this the sum-count method. If the datapoints $x_{l}$ are contained in $[-B,B]$, then the sensitivity of the sum is $B$, and the sensitivity of the count is $1$. Adding noise at these sensitivities yields reasonable utility, since the sum has an expected magnitude much larger than $B$, so the noise does not dominate the value, and upon division, the noise is further reduced.
The sum-count method is typically more privacy-efficient than releasing the mean directly. If we release the mean directly, the sensitivity needs to account for the existence of clusters with a single data point, leading to a sensitivity of $B$ for the mean. Since the magnitude of the mean itself is at most $B$, a sensitivity of $B$ produces a relatively large amount of noise, leading to unusable utility.
We show that by using constrained clustering, we can tighten the sensitivity bound and release the mean directly. This avoids dividing the privacy budget between the sum and count queries and gives a much higher utility than the sum-count method as we show in Section 6.2. By restricting the minimum cluster size to $\eta_{min}$, we provide a tunable bound on the sensitivity of the mean. Since all clusters have at least $\eta_{min}$ data points, the denominator of the centroid is now bounded below by $\eta_{min}$. Therefore, a single data point cannot affect any centroid by more than $\frac{B}{\eta_{min}}$.
(6) $\Delta^{\mu}=\frac{B}{\eta_{min}}$
The centroid-method for computing the DP mean is:
(7) $\bar{\mu_{j}}=\frac{\sum_{x_{l}\in O_{j}}x_{l}}{|O_{j}|}+\gamma^{\mu}\quad\text{where}\quad\gamma^{\mu}\sim Lap\left(\frac{\Delta^{\mu}}{\epsilon^{\mu}}\right)$
We further claim that, by employing constrained clustering, the sum-count method can also be improved. Intuitively, having a minimum cluster size prevents the formation of very small clusters, where the differentially-private noise would dominate the underlying sum and count. In Section 6.2, we show that the sum-count method is improved by using constrained clustering, but not as much as the centroid method.
5.2. Overview of FastLloyd
In this section, we describe our protocol for federated DP $k$-means, FastLloyd. Algorithm 1 overviews the protocol from the perspective of a single client. Following the structure of Lloyd’s algorithm, the following sections describe how we design each of its main steps: initialization, assignment, and update.
5.2.1. Initialization: Sphere Packing
We modify the initialization so that it can be carried out in a federated manner. We employ the sphere packing initialization approach of Su et al. (Su et al., 2016) as it was shown to outperform random initialization. The sphere packing approach is data-independent and thus does not use any privacy budget. The process can be outlined as follows:
  1. (1)
    Initialize a radius parameter $a$.
  2. (2)
    For $j\in\{1,2,...,k\}$, generate a point $\mu_{j}$ such that it is at least of distance $a$ away from the domain boundaries and at least of distance $2a$ away from any previously chosen centroid. If a randomly generated point does not meet this condition, generate another one.
  3. (3)
    If, after repeated attempts, it is not possible to find such a point, decrease the radius $a$ and repeat the process.
The radius $a$ is determined via a binary search to find the maximum $a$ that allows for the generation of $k$ centroids. In Algorithm 1, each client independently calls the Initialization function (Line 3) with the same random seed, which results in each client starting with the same centroids.
5.2.2. Assignment: Constrained $k$-means
To set the minimum cluster ($\eta_{min}$) and maximum cluster ($\eta_{max}$) sizes, we apply a heuristic of setting them as a ratio of the average cluster as follows:
(8) $\eta_{min}=\frac{N}{\alpha_{min}k}\quad\text{and}\quad\eta_{max}=\frac{\alpha_{max}N}{k}$
where $\alpha_{min}$ and $\alpha_{max}$ are hyperparameters that control the minimum and maximum cluster sizes. In Appendix B, we show that $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ provide a good balance between utility and feasibility. We use the open-source implementation from Levy-Kramer (Levy-Kramer, 2018) to enforce the constraints in the assignment step. In Line 6 of Algorithm 1, the client computes their assignments locally using the min-cost flow solver.
Algorithm 1 FastLloyd from $p_{i}$’s perspective
1:Inputs: Local Dataset $D_{i}$.
2:Output: Cluster Centres $\boldsymbol{\mu}^{(T)}$
3:$\boldsymbol{\mu}^{(0)}=$ Initialization($seed$)
4:for $t\in\{1:T\}$ do
5:     // Assignment Step
6:     Compute $\mathbf{O}_{i}^{(t)}=$ MinCostFlowSolve($D_{i},\boldsymbol{\mu}^{(t-1)}$)
7:     // Local Update
8:     for $j\in\{1,\dots,k\}$ do
9:         Compute $\bar{\mu}_{ij}^{(t)}=\frac{\sum_{x_{l}\in O_{ij}^{(t)}}x_{l}}{|O_{ij}^{(t)}|}$
10:     // Global Update
11:     $\widehat{\boldsymbol{\mu}}=$ GlobalMSA($\mathbf{\bar{\mu}}_{i}^{(t)}/M$, $\Delta$, $\epsilon^{\mu},seed,t$)
12:     // Post Process Result
13:     $\boldsymbol{\mu}^{(t)}=$ Fold($\widehat{\boldsymbol{\mu}}$)
5.2.3. Local Update
The client has already locally assigned each of their data points to a cluster, following the constraints, in the previous step (Line 6). Next, they compute the sum divided by the count of each cluster, using their local dataset. We call this the local update, and it follows the classic Lloyd’s algorithm exactly. The output of this step is a matrix (of dimensions $(k,d)$) containing the average of each dimension for each cluster on client $p_{i}$’s data.
5.2.4. Global Update: Masked Secure Aggregation
We modify the update step to release a perturbed version of the cluster centres to each client under the CDP security model. By publishing the previous noisy centroids, the assignment and local centroid computation can be performed locally by each client instead of using a secure computation protocol. In the global update step (Line 11), we privately aggregate (average) the local centroids of all clients, revealing the noisy global centroids for the next round. The clients scale their local centroids down by $M$, before calling the global update so that the sum of the local updates results in the average of the local centroids.
We describe our aggregation protocol in Figure 2. We call this protocol Masked Secure Aggregation (MSA). In MSA, similar to regular secure aggregation protocols (Bonawitz et al., 2017), a service provider S securely aggregates values from $M$ clients, all while oblivious to every client’s contributions. However, in MSA, the server is also oblivious to the result of the aggregation operation; the server only acts as an aggregator (who also adds DP noise) and is oblivious to both the input and output. Each client $p_{i}$ possesses a private value (a matrix $v_{i}\in\mathbf{R}^{k\times d}$, where $\mathbf{R}$ could be $\mathbb{Z}_{2^{64}}$). These clients aim to collectively compute the element-wise sum of their private matrices: $\boldsymbol{v}=\sum_{i=1}^{M}{\frac{v_{i}}{M}}$. In addition to the client’s values, we assume the privacy parameters ($\epsilon,\Delta$) and a shared random $seed$ are also provided.
Clients Send Data
The first step in the protocol is for each of the $M$ clients to generate a random mask set. The random mask set is composed of $M$ random matrices of the same dimensions as the client items: $\{r_{1},r_{2},\dots r_{M}\}$ where each value is sampled uniformly from $\mathbf{R}$. Because all the clients have access to a shared seed and $PRNG$ they can each compute the entire set locally. In Line 2, each client samples the entire set and sums it (in Line 2) to get the global mask matrix $\boldsymbol{r}$ (which will be used to decrypt later).
In Line 2, each private value is converted to a fixed-point format, $\tilde{v_{i}}$, by scaling it up with a power-of-2 scale factor, $\mathbf{2^{\delta}}$. Formally, $\tilde{v_{i}}=\lceil{v_{i}\times\mathbf{2^{\delta}}}\rfloor$. All fractional values are rounded to the nearest representable value in this fixed-point representation. The choice of the scale factor, $\mathbf{2^{\delta}}$, determines the precision of the representation [^1] A larger $\mathbf{2^{\delta}}$ allows for greater precision but also reduces the number of bits available for the integer part of the number, which might cause overflows. A workaround is to increase the bit-width of the operations, which increases the computational load and the communication cost.. In practice, we empirically choose $\delta=16$.
Then, in Line 2, each client encrypts (masks) their input $\tilde{v_{i}}$ by adding their share of the mask $r_{i}$. Finally, each client sends their masked values to the server. This step masks the actual value $v_{i}$ from the server because $r_{i}$ acts as a one-time pad. Note that this masked value is never sent to other clients, as they would be able to unmask it easily.
[…]
Figure 2. Global Masked Secure Aggregation Protocol with $M$ Clients.
Server Aggregates Data
Upon receiving the masked matrices, the server first sums over each client’s contribution in Line 2. This yields the masked global sum $\boldsymbol{v}+\boldsymbol{r{}}$. The server then samples from a Laplace distribution to add noise proportional to $\Delta/\epsilon$ for each entry in the result matrix. This noise must also be converted to fixed-point by computing $\tilde{\gamma}=\lceil{\gamma\times\mathbf{2^{\delta}}}\rfloor$. We show why this preserves differential privacy in Appendix A. Finally, the server adds the noise $\tilde{\gamma}$ to the sum and broadcasts it back to the clients.
Client Unmasks Data
Upon receiving the result back from the server, each client must unmask the result. They do this by simply subtracting the mask $r{}$ that they computed in Line 2. After unmasking, each client must scale down the result to retrieve the correct answer. This is done by dividing the unmasked sum by the scale factor, $\mathbf{2^{\delta}}$, to reverse the initial scaling operation.
5.2.5. Postprocessing
In Algorithm 1, each client locally postprocesses the results (since adding noise to the centroids can result in values outside of the domain). A naive postprocessing strategy is to simply truncate out-of-bounds centroids to the boundary. However, in practice, we find that folding (Holohan et al., 2019) the value (reflecting it over the boundary) gives better utility. More formally, if $x<-B$, we adjust it to $x^{\prime}=-2B-x$, and if $x>B$, we adjust it to $x^{\prime}=2B-x$. This folding process effectively “folds” (mirrors) the value over the nearest boundary, preserving its distance from the boundary. We show in Appendix B that folding improves the utility of the algorithm, especially when the noise is large (small privacy budget).
5.3. Proof of Privacy
FastLloyd allows for a central party to add noise, retaining the utility of the central model of DP. However, since the server only interacts with masked values, and the clients can only observe a differentially private view of the final (noised) centroids, the protocol provides a level of privacy akin to that of the local model of DP. We prove the end-to-end privacy of our algorithm in the following theorem.
Theorem 5.1.
Algorithm 1 ensures $\epsilon(\lambda)$-IND-CDP-MPC in the presence of a semi-honest, polynomial time adversary who controls at most a single party. Where
$\epsilon=Td\epsilon^{\mu}$
Proof.
In order to prove the algorithm satisfies Definition 3.1, we need to consider the view from each party. We begin with the view of the server. First, let us assume that each client samples the random mask $r_{i}$ from a random oracle over $\mathbf{R}$. Under this assumption, the $Enc(\cdot)$ function satisfies information-theoretic security as it is a one-time pad using $r_{i}$ as the pad. In practice, we implement the sampling of $r_{i}$ using a PRNG and thus reduce from information-theoretic to computational security with a negligible term $negl(\lambda)$. Therefore, the view of the server satisfies $\epsilon(\lambda)$-IND-CDP-MPC with $\epsilon=0$.
In regards to the clients, we assume no client colludes with any other client. Then, a client’s view consists of their own dataset $D_{A}$, and the output from each iteration of the protocol. We note that the initialization is data-independent and thus is indistinguishable between neighbouring datasets. Each iteration consists of the assignment and updating of the cluster centroids. The assignment step is simply postprocessing on the previous iteration (or random initialization) and thus incurs no additional privacy cost. The updating of the cluster centroids applies the Laplace mechanism to each dimension separately, with the sensitivity parameter set following Eqn 6 and $\epsilon^{\mu}$ as the privacy budget. Using naive, sequential composition over the dimensions and the iterations, the client’s view is $Td\epsilon^{\mu}$-IND-CDP-MPC, with no computational assumption (as we use the information-theoretic DP properties of the Laplace mechanism and composition (Dwork and Roth, 2014)). Taking the worst case over the clients and the server, the result follows. ∎
5.4. Error Analysis
To analyze the error of our approach, we follow a similar analysis to Su et al. (Su et al., 2016). We analyze the mean-squared error between the true centroid ($\mu_{j}$) and the differentially-private centroid ($\bar{\mu_{j}}$) for one iteration across all dimensions. Substituting the definition of the centroid method (Eqn 7), we get
$\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=\mathsf{E}\left[\sum_{h=1}^{d}((\mu_{j}[h]+\gamma^{\mu})-\mu_{j}[h])^{2}\right]=$
$=\sum_{h=1}^{d}\mathsf{E}\left[(\gamma^{\mu})^{2}\right]=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right).$
Where the last line holds because $\gamma^{\mu}$ is an independent Laplacian variable with zero mean and so
$\mathsf{E}[(\gamma^{\mu})^{2}]=\mathsf{Var}\left(\gamma^{\mu}\right)-(\mathsf{E}[\gamma^{\mu}])^{2}=\mathsf{Var}\left(\gamma^{\mu}\right).$
Plugging in the sensitivity of the mean (Eqn 6) and the privacy budget gives
(9) $\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right)=\frac{2dB^{2}}{(\epsilon^{\mu}\eta_{min})^{2}}$
(10) $=\frac{2d^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}=\mathcal{O}\left(\frac{d^{3}T^{2}}{\epsilon^{2}\eta_{min}^{2}}\right).$
This shows that the distortion in the centroid is proportional to $d^{3}T^{2}$ and inversely proportional to $\epsilon^{2}\eta_{min}^{2}$.
Remark.
This error is similar to what Su et al. (Su et al., 2016) observed for the sum-count method with the following differences: (1) $\eta_{min}$replaces the assumed expected cluster size of $\frac{N}{k}$. (2) The error is independent of the data or the actual clustering result; it depends completely on the parameters set before running the algorithm. (3) This is a worst-case analysis, and the actual error might be lower.
Adding across all clusters gives
(11) $\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}.$
Using this analysis, we can determine the optimal number of iterations, following a similar approach as Su et al. (Su et al., 2016). First, we calculate a minimum privacy budget per iteration ($\epsilon^{(m)}$) so that the mean-squared error is within a certain bound. Secondly, we divide the available budget ($\epsilon$) by $\epsilon^{(m)}$ in order to determine the number of iterations $T$. Thus, setting $T=1$ and $\epsilon=\epsilon^{(m)}$ gives us
$\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}B^{2}}{(\epsilon^{(m)}\eta_{min})^{2}}\leq 0.004$
which implies
$\sqrt{\frac{500kd^{3}B^{2}}{\eta_{min}^{2}}}\leq\epsilon^{(m)}.$
However, Su et al. note that, in practice, there is no need to go beyond seven iterations and that at least two iterations are needed to gain useful results. Therefore, we truncate the number of iterations to be in the range $[2,7]$.
6. Evaluation
In this section, we provide an extensive evaluation of our work in terms of the following questions:
  1. Q1
    How does the utility of FastLloyd  compare with state-of-the-art in central model (Su et al., 2016) for DP $k$-means?
  2. Q2
    How does constrained clustering affect the utility of the sum-count method?
  3. Q3
    How does the utility of FastLloyd scale with varying the number of dimensions and number of clusters?
  4. Q4
    How does the runtime, communication, and number of rounds for FastLloyd compare with the state-of-the-art in federated $k$-means using secure computation (Mohassel et al., 2020)?
  5. Q5
    How does the runtime and communication of FastLloyd  scale with varying the number of dimensions, clusters, and data points?
6.1. Experimental Setup
(a) S1
(b) Lsun
(c) Birch2
(d) House
(e) Iris
(f) Adult
Figure 3. Utility evaluation of FastLloyd vs. SuLloyd on real datasets (20 runs).
Implementaion
The experiments were carried out on a Macbook Pro M2 Max, with a 30-core CPU, a 38-core GPU, and 64 GB of RAM. We use the Open MPI (Gabriel et al., 2004) (Open Source High-Performance Computing) library to implement multiparty communication through message-passing. All experiments were conducted in the multiparty framework, with datasets being randomly split between parties.
Following Mohassel et al. (Mohassel et al., 2020), we only evaluate in the LAN setting, and observe that WAN runtimes can be computed with the linear cost model (adding a fixed latency for each round). We simulate the network latency by introducing a custom (sleep) delay of 0.5ms for every “send” operation. We make our code publicly available to reproduce the results222 https://github.com/D-Diaa/privateclustering.git. All results were averaged over 20 runs and the 95% mean confidence interval is supplied where appropriate.
Baselines
To evaluate the clustering quality of FastLloyd, we compare it against the state-of-the-art in central DP for $k$-means clustering by Su et al.  (Su et al., 2016), denoted as SuLloyd. As established in Section 4, SuLloyd is the most accurate DP $k$-means clustering algorithm to date. We implement a federated version of SuLloyd in our framework for fair evaluation. To study the effect of constrained clustering on the sum-count method used in SuLloyd, we also extend the federated implementation of SuLloyd to use the same constraints as FastLloyd, denoted as CSuLloyd. Finally, we give a non-private baseline of the original Lloyd’s algorithm (Lloyd, 1982), implemented into our federated framework (with sphere packing initialization) and denoted as Lloyd.
To evaluate the computational and communicational costs in the secure federated setting, we compare against the corresponding state-of-the-art by Mohassel et. al (Mohassel et al., 2020), denoted as MohLloyd. According to the SoK of Hedge et al. (Hegde et al., [n. d.]), MohLloyd is the most efficient secure $k$-means clustering algorithm to date. For MohLloyd, we evaluate against the reported values in the original paper (Mohassel et al., 2020).
Parameters
In our experiments, we use the following default parameters, unless otherwise specified. The number of clients is set to two, and the privacy budget $\epsilon=1$. For the MSA protocol (Figure 2), the underlying ring $\mathbf{R}$ is $\mathcal{Z}_{2^{32}}$ and the values are all transformed to fixed-point under precision factor $\delta=16$. For any dataset of size $N{}$, to be clustered into $k{}$ clusters, we fix $\eta_{min}=\frac{N{}}{1.25k{}}$ and $\eta_{max}=\frac{1.25N{}}{k{}}$ (empirically chosen as per Appendix B).
Datasets
We selected a wide range of commonly used datasets to evaluate our approach, we follow Su et al. (Su et al., 2016) and Mohassel et al. (Mohassel et al., 2020) to maximize comparability. An exception is the Birch2 (Zhang et al., 1997) dataset, which is a synthetic dataset comprised of 100 Gaussian clusters forming a sine wave. This dataset initially contains 100K data points, however, a random sample of 25K points is used. To the best of our knowledge, we are the first federated privacy-preserving work to evaluate using a dataset of this size (in terms of $N$ and $k$).
Additionally, we created synthetic datasets to evaluate the scalability of our approach. Similar to Su et al. (Su et al., 2016), we use the R package (Qiu and Joe, 2023), clusterGeneration. The package is designed for generating cluster datasets with a given degree of separation ($\in[-1,1]$) between neighbouring clusters. For the clustering quality evaluation, we generate two sets of datasets: one with overlapping cluster structure (SynthOV, separation = 0.21) and one with well-separated cluster structure (SynthSEP, separation = 0.70). Each set varies $k$ and $d$ from $2$ to $20$ in steps of $2$, resulting in 100 variations. For every variation, we generate three different datasets with different seeds. For all synthetic datasets, we fix an average cluster size of $C_{avg}=\frac{10K}{k}$ and randomly sample cluster sizes to be in the range $[0.75\cdot C_{avg},1.25\cdot C_{avg}]$, resulting in a dataset that has a total size in the range $[7.5K,12.5K]$. For our timing evaluation, we generate similar synthetic datasets, TimeSynth, similar to those in Mohassel et al.’s work (Mohassel et al., 2020), enforcing an exact cluster size of $C_{avg}=\frac{N}{k}$. We vary $N$ between 10K and 100K, $d$ between two and five and $k$ between two and five. A summary of these datasets is presented in Table 2. As is common in the DP literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), we assume all datasets are normalized in the range $[-1,1]$ as pre-processing.
Table 2. Summary of Datasets Used in Evaluation
Dataset $N$ $d$ $k$
Iris (Fisher, 1988) 150 4 3
LSun (Ultsch, 2005) 400 2 3
S1 (Fränti and Virmajoki, 2006) 5000 2 15
House (Fränti and Sieranoja, 2018) 1837 3 3
Adult (Becker and Kohavi, 1996) 48842 6 3
Birch2 (Zhang et al., 1997) 25000 2 100
SynthOV 7.5K ~ 12.5K 2-20 2-20
SynthSEP 7.5K ~ 12.5K 2-20 2-20
TimeSynth 10K, 100K 2,5 2,5
6.2. Utility Evaluation
In this section, we focus on evaluating the utility of FastLloyd compared to the state-of-the-art work of Su et al. (Su et al., 2016). We begin by describing the metrics used for this evaluation.
Metrics
Following similar literature (Su et al., 2016), we use the Normalized Intra-cluster Variance (NICV) metric to evaluate clustering quality. In short, NICV is the $k$-means objective function (1) divided by the number of points in the dataset.
$NICV(O)=\frac{1}{N}\cdot\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
We also introduce another metric, Empty Clusters, that measures the ratio of empty clusters to $k$. The objective of $k$-means clustering is to produce exactly $k$ clusters such that the WCSS (NICV) is minimized. However, existing metrics/objectives do not capture the requirement of having exactly $k$ clusters. This becomes problematic in differentially private settings, where adding noise can isolate some centroids from the data completely, resulting in a suboptimal solution that uses less than $k$ clusters. We introduce this metric to precisely quantify this effect as follows:
Empty $\texttt{Empty Clusters}(O)=\frac{\sum_{j=1}^{k}{I(|O_{j}|=0)}}{k}$
6.2.1. Real World Datasets
To answer Q1 and Q2, we evaluate FastLloyd, CSuLloyd, and SuLloyd on a variety of datasets. Figure 3 shows the NICV against various privacy budgets ($\epsilon$) from 0.01 to 2.0 on six datasets. The shading shows the 95% confidence interval over the 20 experiments.
First, we consider Q1. We observe that FastLloyd outperforms SuLloyd on all datasets, especially for lower values of epsilon. On a larger dataset (birch2), Figure 3(c) shows a statistically significant improvement (the confidence intervals do not overlap). FastLloyd even outperforms the non-private baseline of Lloyd at $\epsilon=2$. For higher dimensional datasets, a substantial performance gap is evident; Figure 3(f) shows that even in high-privacy settings (with $\epsilon=0.01$), FastLloyd significantly outperforms the non-private baseline of Lloyd. To answer Q2, we observe that for all datasets CSuLloyd  outperforms SuLloyd. However, in all cases, FastLloyd outperforms SuLloyd (except in Figure 3(f) , where we tie).
To evaluate the Empty Clusters metric, we use the birch2 dataset with a relatively high cluster count of 100. Figure 4 shows that even the non-private Lloyd loses around $15\%$ of clusters, with SuLloyd reaching up to more than $40\%$ in high-privacy settings (with $\epsilon=0.01$). We observe that FastLloyd eliminates this problem completely for $\epsilon>0.75$ and significantly diminishes it for smaller privacy budgets.
Figure 4. Birch2 Empty Clusters
(a) SuLloyd
(b) FastLloyd
(c) Improvement Percentage
Figure 5. NICV Utility comparison over SynthOV dataset. Averaged over 60 runs (20 on each of three random datasets).
(a) SuLloyd
(b) FastLloyd
(c) Improvement Percentage
Figure 6. NICV Utility comparison over SynthSEP dataset. Averaged over 60 runs (20 on each of three random datasets).
6.2.2. Synthetic Datasets
To answer Q3, we follow an approach similar to Su et al. (Su et al., 2016) and evaluate FastLloyd and SuLloyd on SynthSEP and SynthOV, varying both $k$ and $d$. Figures 5 and 6 show heatmaps of NICV values for both approaches when $\epsilon=1$. To clarify the difference between the protocols, Figures 5(c) and 6(c) show the percentage of reduction in NICV of FastLloyd over that of SuLloyd. For the overlapping dataset (SynthOV), FastLloyd  outperforms SuLloyd in all cases, with the improvement being more significant in higher cluster counts and low dimensions, reducing NICV by more than 40%. The gap in performance becomes more pronounced in the separated dataset (SynthSEP), with FastLloyd  significantly outperforming SuLloyd in all cases, reducing the NICV error by up to 89% in some cases, and by more than 78% even in the highest number of clusters and dimensions. These observations highlight the better scalability and utility of FastLloyd over SuLloyd in high-dimensional and high-cluster-count scenarios. Since SuLloyd was found to outperform all other approaches for $d{}>3$ (Su et al., 2016), we claim that FastLloyd is state-of-the-art in DP $k$-means for $d{}>3$. FastLloyd also fixes the scalability issue that Su et al.’s work (Su et al., 2016) observed in high $k$.
6.3. Runtime Evaluation
Dataset Parameters Runtime (ms) Comm (bytes)
$N$ $k$ $d$ FastLloyd MohLloyd (Mohassel et al., 2020) Speedup FastLloyd MohLloyd (Mohassel et al., 2020) Reduction
TimeSynth 10K 2 2 $3.23\pm 0.23$ 10500 3250×3250$\times$3250 × 128 2.557e8 1,998,000×1,998,000$\times$1 , 998 , 000 ×
2 5 $8.61\pm 1.81$ 320 - -
5 2 $18.02\pm 3.59$ 34050 1890×1890$\times$1890 × 320 9.746e8 3,045,000×3,045,000$\times$3 , 045 , 000 ×
5 5 $19.83\pm 2.01$ 800 - -
TimeSynth 100K 2 2 $85.57\pm 2.42$ 105120 1228×1228$\times$1228 × 128 2.467e9 19,277,000×19,277,000$\times$19 , 277 , 000 ×
2 5 $313.47\pm 6.79$ 320 - -
5 2 $423.1\pm 6.88$ 347250 820×820$\times$820 × 320 9.535e9 29,797,000×29,797,000$\times$29 , 797 , 000 ×
5 5 $412.89\pm 5.84$ 800 - -
LSun 400 3 2 $1.81\pm 0.1$ 1481 1156×1156$\times$1156 × 192 - -
S1 5K 15 2 $24.76\pm 0.3$ 49087 1982×1982$\times$1982 × 960 - -
Table 3. Overhead comparison per iteration against MohLloyd (Mohassel et al., 2020) for two clients ($20$ runs with mean and $95\%$ confidence reported for FastLloyd)
To answer questions Q4 and Q5, we compare FastLloyd with the MohLloyd protocol of Mohassel et al. (Mohassel et al., 2020) on a variety of common datasets in terms of runtime and communication size (per iteration). Table 3 shows the summary of the evaluation with two clients (the setting considered in Mohassel et al.’s work (Mohassel et al., 2020)). Values for MohLloyd are taken as reported in the paper by Mohassel et al. (Mohassel et al., 2020), noting that the setup is similar to the one used in our evaluation, and the gap in performance is more than what could be accounted for by different setups. In terms of runtime, MohLloyd executes in the order of minutes, while FastLloyd executes in a fraction of a second, offering four orders of magnitude speedup. In terms of communication size, MohLloyd requires communicating gigabytes of data per iteration, while FastLloyd needs a fraction of a kilobyte, offering up to seven orders of magnitude reduction in size. In terms of communication rounds, MohLloyd requires $\Theta(\lceil\log{k{}}\rceil)$ communication rounds per iteration (Hegde et al., [n. d.]), while FastLloyd only requires one. Since MohLloyd was found to be the state-of-the-art in secure federated $k$-means (Hegde et al., [n. d.]), we conclude that we advance the state-of-the-art while offering output privacy (which would only further slow down Mohassel et al. (Mohassel et al., 2020)).
7. Extensions
7.1. Distance-Constrained $k$-means
(a) S1 NICV
(b) Birch2 NICV
(c) Birch2 Empty Clusters
Figure 7. The effect of the distance constraint (dashed lines) on the clustering quality of FastLloyd and CSuLloyd
[…]
Figure 8. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$, $\boldsymbol{\eta}^{max}$ and $\eta^{dist}$.
We introduce a new extension of the constrained clustering problem that further tunes and improves the utility of DP clustering. This new problem includes a constraint on the maximum radius ($\eta_{d}$) of every cluster, such that the objective function is now subject to the constraints:
(12) $\eta_{min}\leq|O_{j}|\leq\eta_{max}\quad\text{and}\quad||x_{l}-\mu_{j}||\leq\eta_{d}\quad\forall x_{l}\in O_{j},\forall j$
In lieu of previous work, we extend the network flow formulation to solve our new constrained clustering problem; an example is shown in Figure 8. In short, we modify the capacity of edges connecting the data points to the centroids to be $1$ if the distance is less than or equal to $\eta_{d}$, and 00 otherwise, essentially removing the edge. This prevents data points from being assigned to centroids that are further than $\eta_{d}$ away, even if it means that they do not get assigned to any cluster.
Intuitively, this constraint has several benefits. First, it provides a form of resistance to outliers, as data points that are far away from any centroid will not (and in some cases, should not) be assigned to any cluster. Second, it allows for more fine-grained control over the clusters, as the maximum distance constraint can be used to ensure that the clusters are compact. Most importantly, the maximum distance constraint helps to tighten the sensitivity bound of the centroids even further.
Theorem 7.1.
If constraints (12) are satisfied, then:
(13) $\Delta^{\mu}\leq\frac{2\eta_{d}}{\eta_{min}}$
Proof.
Since the update phase treats every dimension separately (see Section 5.3, we can consider only one dimension. Given a previous centroid $\mu_{t-1}$ (that might be randomly initialized), we need to bound the maximum change in the new centroid $\mu_{t}$, when changing the neighbouring dataset. Specifically, we bound the following
(14) $\Delta^{\mu}=\max\limits_{D\sim D^{\prime}\in\mathcal{D}}||\mu_{t}(D,\mu_{t-1})-\mu_{t}(D^{\prime},\mu_{t-1})||_{1}.$
In the unbounded model, $D\sim D^{\prime}$ means that the datasets are identical except for one data point, $x^{\prime}$, which is in $D^{\prime}$ but not in $D$.
If $x^{\prime}$ is not within $\eta_{d}$ of $\mu_{t-1}$, it will not be assigned to the new cluster, and the neighbouring centroids will be the same, completing the proof. Therefore, we only need to consider the case where $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$. In this case, we have two neighbouring assignments, $O_{t}$ and $O_{t}^{\prime}=O_{t}\cup\{x^{\prime}\}$, and the corresponding centroids will be:
$\mu_{t}$ $=$ $\frac{\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|}$
$\mu_{t}^{\prime}$ $=$ $\frac{x^{\prime}+\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|+1}$
Now, letting $\Sigma=\sum_{x_{l}\in O_{t}}x_{l}$, and $C=|O_{t}|$ to simplify the notation, we bound the difference between the neighbouring centroids as follows:
(15) $||\mu_{t}-\mu_{t}^{\prime}||_{1}$ $=\left|\left|\frac{\Sigma}{C}-\frac{\Sigma+x^{\prime}}{C+1}\right|\right|_{1}=\left|\left|\frac{\Sigma-Cx_{l}}{C(C+1)}\right|\right|_{1}$
(16) $\leq\left|\left|\frac{\Sigma-C(\mu_{t}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
(17) $=\left|\left|\frac{\Sigma-C(\frac{\Sigma}{C}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
(18) $=\left|\left|\frac{\mp 2\eta_{d}C}{C(C+1)}\right|\right|_{1}\leq\frac{2\eta_{d}}{C}\leq\frac{2\eta_{d}}{\eta_{min}}$
where (18) follows from the fact that $C\geq\eta_{min}$ and (16) follows from the fact that $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$, and therefore by the triangle inequality:
$||x_{l}-x^{\prime}||_{2}\leq||x_{l}-\mu_{t-1}||_{2}+||\mu_{t-1}-x^{\prime}||_{2}\leq 2\eta_{d}\quad\forall x_{l}\in O_{t}$
which implies that $x^{\prime}$ is within $2\eta_{d}$ of any point in $O_{t}$, namely their average $\mu_{t}$. Since the $\ell_{1}$ norm over one dimension ($\ell_{\infty}$) is always less than or equal to the $\ell_{2}$ norm, we have:
$||x^{\prime}-\mu_{t}||_{1}\leq||x^{\prime}-\mu_{t}||_{2}\leq 2\eta_{d}$
implying that, in this dimension, $x^{\prime}=\mu_{t}\pm 2\eta_{d}$ in the worst case. ∎
We heuristically set this parameter as follows (for 2 dimensions):
(19) $\eta_{d}=\sqrt{\frac{B^{2}}{2t}}$
The intuition behind this choice is that it allows a larger distance ($\eta_{d}\approx 0.7B$) at the beginning of the algorithm (to allow the initial centroids to align with the data) and then gradually reduces the distance, to reduce the noise and fine-tune the clusters. In particular, we propose that $\eta_{d}$ should be a function of $k$ and $d$, as the higher number of clusters implies the more compact they are expected to be, and the higher number of dimensions implies that the data points are likely to be more spread out.
A preliminary evaluation using the S1 and Birch2 datasets shows that this constraint can significantly improve the utility of the algorithm. Figure 7 shows this improvement on both the constrained sum-count method CSuLloyd and the constrained centroid method FastLloyd, where we observe a $30\sim 40\%$ reduction in NICV in small $\epsilon$. The trade-off with this maximum radius constraint is that the network flow becomes harder to satisfy. We find that for other datasets, such as Adult, the min-cost flow solver does not always converge. We leave further exploration of this constraint to future work.
7.2. Serverless Extensions
In this section, we discuss alternatives to using the service provider. A trivial alternative is that each client could locally noise their centroids before sending them to another client (or set of clients) to aggregate. The downside to this approach is that the error will compound with the number of clients and decrease utility. An alternative is for all or a subset of the clients to take the role of a computation node in an MPC protocol. In this case, the centroids would be secret shared and aggregated in MPC, then Laplace noise would need to be added similar to FastLloyd. The noise addition can be implemented by each computation node sampling noise locally and adding it to their shares. Depending on how many nodes are assumed to be honest, this will necessarily add more noise than FastLloyd to ensure privacy in the presence of one or more nodes colluding. Alternatively, the computation nodes can jointly sample the noise using the protocol of Wu et al. (Wu et al., 2016). The disadvantage of this approach is that it will add a large computational overhead which will make the runtime closer to that of Mohassel et al. (Mohassel et al., 2020).
8. Conclusion
In this work, we designed FastLloyd, a new private protocol for federated $k$-means. We have shown that FastLloyd is secure in the computational model of DP and analyzed its utility. Compared to state-of-the-art solutions in the central model of DP, FastLloyd results in higher utility across a wide range of real datasets and scales effectively to larger dimensions and number of clusters. FastLloyd also achieves four orders of magnitude faster runtime than the state-of-the-art in secure federated $k$-means across a variety of problem sizes. In summary, we provide an efficient, private, and accurate solution to the horizontally federated $k$-means problem.
Acknowledgements.
We gratefully acknowledge the support of the National Research Council Canada, the University of Waterloo, the Government of Ontario, and the NSERC Postgraduate Scholarship-Doctoral program for funding this research.
Appendix A Proof of Quantized DP Noise
We argue why quantizing the DP noise is acceptable in Figure 2. In essence, because the sensitive data is already quantized, adding quantized noise to it is equivalent to adding the noise first and then performing the quantization, which aligns with the postprocessing lemma in differential privacy.
This can be intuitively understood by noting that:
  • The data before noise addition is already quantized.
  • The sum of a quantized value and a non-quantized value will only have information of the non-quantized variable in the less significant bits (i.e., those lost to quantization).
  • Hence, quantizing the sum is equivalent to dropping this lower bit information, which is similar to quantizing the second variable prior to addition.
We show this formally in the following Theorem.
Theorem A.1 (Quantization and Differential Privacy).
Quantizing Laplace noise at the same level of quantization as that used in the Masked Secure Aggregation (MSA) protocol does not violate the privacy guarantees offered by differential privacy.
Proof.
We begin by observing that any value, say $v$, can be split into two parts upon quantization: the integral part, $\bar{v}$, and the fractional part, $\hat{v}$, such that $\bar{v}\times\mathbf{2^{\delta}}$ is the (rounded) integral part of $v\times\mathbf{2^{\delta}}$ (i.e. it is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$), and $\hat{v}$ denotes the fractional part that gets lost due to quantization. This notation can similarly be applied to the noise variable, $\eta$.
In our context, the sensitive data we wish to protect is $\bar{v}$ since the entire protocol operates in a quantized environment, i.e., $v$ is never transmitted. We aim to demonstrate the following equivalence:
$\tilde{v}+\tilde{\eta}=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta\times\mathbf{2^{\delta}}}\rfloor=\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor$
The right-hand side represents quantization applied as postprocessing to the sum of the sensitive data and the noise, which adheres to the rules of differential privacy.
After expanding the right-hand side, we obtain:
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{\bar{v}\times\mathbf{2^{\delta}}+\bar{\eta}\times\mathbf{2^{\delta}}+\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
We can take out $\bar{v}\times\mathbf{2^{\delta}}$ and $\bar{\eta{}}\times\mathbf{2^{\delta}}$ from the rounding operator because they are exact integers (indeed $\bar{v}\times\mathbf{2^{\delta}}$ is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$). We then obtain:
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor+\lceil{\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
However, $\hat{\eta}\times\mathbf{2^{\delta}}$ will be lost due to quantization, hence:
$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor=\tilde{v}+\tilde{\eta}$
Appendix B Ablation of Parameters
Figure 9. Ablation over different selections of parameters, $\epsilon=1.0$
Figure 10. Ablation over different selections of $\alpha$, $\epsilon=0.1$
In this section, we empirically justify our choice of hyperparameters: $\eta_{min}$, $\eta_{max}$ and the postprocessing method. We choose these parameters to minimize the Normalized Intra-Cluster Variance (NICV) metric over a set of synthetic datasets. The synthetic datasets were generated using the clusterGeneration package (Qiu and Joe, 2023) varying the number of dimensions ($d$) and clusters ($k$) as $(k,d)\in\{2,4,8,16\}^{2}$, resulting in 16 variations. For each variation, we generate 3 different datasets with the degrees of separation $\{0.25,0.5,0.75\}$. The average cluster size is fixed to $C_{avg}=\frac{2048}{k}$ and the cluster sizes are randomly sampled in the range $[0.70\cdot C_{avg},1.30\cdot C_{avg}]$, leading to a total dataset size in the range $[1433,2662]$.
All the ablation experiments were done on two privacy budget settings: $\epsilon=1.0$ and $\epsilon=0.1$. The results are shown in Figures 9 and 10, respectively. For the min and max cluster experiments, we plot the average of all the NICV values for each $k$ as a separate line and the average of all datasets over all variations as the black dotted line, with the 95% confidence interval shaded around it. We observe that generally, a smaller $\alpha_{min}$ and $\alpha_{max}$ lead to better NICV values, especially for datasets with a high number of clusters. This is expected because a smaller $\alpha_{min}$ gives a tighter bound on the sensitivity and so improves the utility. A smaller $\alpha_{max}$ also imposes a tighter maximum cluster size constraint, which intuitively helps with the convergence of the algorithm. However, the tighter the constraints, the harder it is for them to be satisfied. As we observe in the plots, most datasets with $\alpha_{min}$ or $\alpha_{max}$ set to lower than $1.25$ do not have a value because the min-cost flow solver cannot find a valid assignment. Therefore, we choose $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ as a balance between utility and feasibility. We also observe that the folding method is better in the low privacy setting ($\epsilon=0.1$), and it does not make a significant difference in the high privacy setting ($\epsilon=1.0$). Therefore, we choose the folding method as the default postprocessing method.

### v2

@@
 gamecenter,boxed
-Abstract.
+Abstract
 We study the problem of privacy-preserving $k$-means clustering in the horizontally federated setting.
-Existing federated approaches using secure computation, suffer from substantial overheads and do not offer output privacy.
-At the same time, differentially private (DP) $k$-means algorithms assume a trusted central curator and do not extend to federated settings.
-Naively combining the secure and DP solutions results in a protocol with impractical overhead.
+Existing federated approaches using secure computation suffer from substantial overheads and do not offer output privacy.
+At the same time, differentially private (DP) $k$-means algorithms either assume a trusted central curator or significantly degrade utility by adding noise in the local DP model.
+Naively combining the secure and central DP solutions results in a protocol with impractical overhead.
 Instead, our work provides enhancements to both the DP and secure computation components, resulting in a design that is faster, more private, and more accurate than previous work.
-By utilizing the computational DP model, we design a lightweight, secure aggregation-based approach that achieves four orders of magnitude speed-up over state-of-the-art related work.
-Furthermore, we not only maintain the utility of the state-of-the-art in the central model of DP, but we improve the utility further by taking advantage of constrained clustering techniques.
-Differential Privacy, Federated Learning, $k$-Means Clustering, Secure Computation
-ccs: Computing methodologies Cluster analysisccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Data anonymization and sanitization
-1. Introduction
+By utilizing the computational DP model, we design a lightweight, secure aggregation-based approach that achieves five orders of magnitude speed-up over state-of-the-art related work.
+Furthermore, we not only maintain the utility of the state-of-the-art in the central model of DP, but we improve the utility further by designing a new DP clustering mechanism.
+1 Introduction
 Unsupervised learning allows data analysts to extract meaningful patterns from raw data that may be difficult to label.
 The canonical example of unsupervised learning is the Euclidean $k$-means clustering problem, where data is grouped into clusters with similar features.
-Clustering has a plethora of important applications in recommendation systems, fraud detection, and healthcare analytics.
+Clustering has a plethora of important applications in recommendation systems, fraud detection, and healthcare analytics [30].
 In such applications, the dataset often contains sensitive data, which necessitates the use of privacy-preserving techniques.
-Combining data from multiple parties in the federated scenario yields more robust insights about the global population at the cost of further exasperating the privacy risks.
-To counteract the privacy risks, a number of works focus on solving the federated $k$-means problem using secure computation (Vaidya and Clifton, 2003; Jha et al., 2005; Jagannathan and Wright, 2005; Jagannathan et al., 2010; Gheid and Challal, 2016; Bunn and Ostrovsky, 2007; Jäschke and Armknecht, 2018; Mohassel et al., 2020; Patel et al., 2012; Yuan and Tian, 2019; Rao et al., 2015; Xiaoyan et al., 2015; Silva and Bellala, 2017; Jiang et al., 2020).
+Combining databases horizontally split across multiple parties in the federated scenario yields more robust insights about the global population at the cost of further exasperating the privacy risks.
+To counteract the privacy risks, a number of works focus on solving the federated $k$-means problem using secure multi-party computation [66, 37, 35, 34, 29, 11, 36, 50, 56, 70, 59, 69, 60, 38].
 Secure computation enables analysts to solve the $k$-means problem while keeping the sensitive data encrypted.
@@
 This gives a false sense of privacy; although the input and intermediate computations all remain private, no effort is made to protect the privacy of the output.
-It is well known in the literature that publishing exact statistics can leak significant information about the input data and allow attacks such as dataset reconstruction (Dwork et al., 2017).
+It is well known in the literature that publishing exact statistics can leak significant information about the input data and allow attacks such as dataset reconstruction [19].
 This level of privacy does not justify the significant runtime overhead incurred by solely using secure computation.
-A separate line of research considers using differential privacy (DP) to protect the privacy of the output of $k$-means (Blum et al., 2005; Dwork, 2011; Su et al., 2016; Mohan et al., 2012; Zhang et al., 2013; Balcan et al., 2017; Lu and Shen, [n.
-d.]; Park et al., 2017; Ni et al., 2021).
-Differential privacy guarantees that the clustering’s output will be approximately the same regardless of any individual user’s participation.
-However, existing methods do not consider the federated setting and assume the existence of a trusted central party who runs the $k$-means algorithm and sanitizes the results with noise before publishing.
-To summarize, no prior work offers a solution to protect both the input and output privacy of the federated $k$-means problem.
-A straw man solution to our problem would be to combine the current state-of-the-art solutions in secure federated $k$-means and DP $k$-means.
+A separate line of research considers using differential privacy (DP) to protect the data.
+Differential privacy guarantees that the output (the cluster centroids) will be approximately the same regardless of any individual user’s participation.
+Differential privacy can be applied in the central, local, or shuffle models.
+In the central model, a trusted aggregator receives the input as plaintext (unprotected) and randomizes the cluster centroids to ensure the output satisfies DP.
+Several works [9, 16, 64, 49, 71, 3, 45, 55, 52] have considered the central setting.
+In the local model, instead of using a trusted aggregator, each party perturbs their data locally before sending it to the aggregator.
+While this removes the need for a trusted aggregator (similar to secure computation), the utility at high privacy levels is low.
+Specifically, the utility is asymptotically worse than the central model by a factor of $\sqrt{n}$, where $n$ is the number of data points [12].
+Several works [54, 62, 13] have considered the local model.
+The shuffle model is a hybrid between the local and central models, with two mutually distrustful parties: a shuffler and an aggregator.
+Despite improving over the local model, the shuffle model still incurs a $6]{n}$ degradation in utility [5].
+To summarize, no prior work offers a solution that protects both the input and output privacy of the federated $k$-means problem and provides a good utility vs.
+privacy trade-off.
+A straw man solution to our problem would be to combine the current state-of-the-art solutions in federated $k$-means using secure computation and $k$-means in the central DP model.
 However, this solution would be unacceptable as the runtime of current federated $k$-means approaches using secure computation is prohibitively large.
 The state-of-the-art approach of Mohassel et al.
-(Mohassel et al., 2020) incurs runtime overheads on the order of tens of minutes.
+[50] incurs runtime overheads on the order of tens of minutes.
 Naively adding differentially private noise to this computation will further increase this runtime.
-Our work focuses on designing an efficient, DP $k$-means in the federated setting that protects the privacy of the input, output, and intermediate computations.
-Following related work in federated $k$-means (Mohassel et al., 2020; Jiang et al., 2020), we use Lloyd’s algorithm as the foundation of our protocol.
-However, we design a new variant of DP-Lloyd (Blum et al., 2005; Su et al., 2016) that noises the centroids directly.
-Our algorithm improves the clustering utility over the state-of-the-art DP algorithm of Su et al.
-(Su et al., 2016) by adding constraints to the $k$-means objective (Bennett et al., 2000).
-Specifically, by adding minimum and maximum constraints to the cluster size, we achieve a non-trivial bound on the sensitivity of the centroids.
-This enables our algorithm to aggregate at the centroid level and apply less DP noise than prior work that aggregates the sums and counts separately (Su et al., 2016).
-To tailor our DP-Lloyd algorithm to secure computation, we design a protocol that publishes the centroids in each step.
+Our work focuses on designing an efficient DP $k$-means in the federated setting (using secure computation) that protects the privacy of the input, output, and intermediate computations.
+Following related work in federated $k$-means [50, 38], we use Lloyd’s algorithm as the foundation of our protocol.
+However, we design a new variant of DP-Lloyd [9, 64] that achieves a tighter bound on the sensitivity.
+Our algorithm significantly improves the clustering utility over the state-of-the-art DP algorithm of Su et al.
+[64], especially in higher dimensions and number of clusters.
+Our algorithm is of independent interest in the central model of DP.
+First, we enforce a bound on the radius of each cluster in the assignment step of Lloyd’s algorithm.
+We then modify Lloyd’s algorithm to compute updates relative to the previous centroid.
+Our relative updates have a sensitivity proportional to the bounded cluster radius rather than the domain size used in previous work [9, 64].
+We find that our radius-based sensitivity bound is best suited to the use of the analytic Gaussian mechanism [6, 15], which further improves utility.
+Finally, we utilize additional post-processing steps based on the radius and the domain to improve the algorithm further.
+Together, these components lead to an improvement of up to 88% in utility (reduction in clustering error).
+To tailor our DP-Lloyd algorithm to secure computation, we design a protocol that publishes intermediate computations in each iteration.
 This allows operations that would be expensive in secure computation (such as assignment and division) to be conducted in plain text.
-However, unlike prior work (Gheid and Challal, 2016; Jiang et al., 2020), we ensure that the intermediate centroids are protected by DP guarantees.
-We prove that the protocol is secure in the computational DP definition (Mironov et al., 2009), which allows DP-bounded leakage during the computation.
+However, unlike prior work [29, 38], we ensure that the intermediate computations are protected by DP guarantees.
+We prove that the protocol is secure in the computational DP definition [48], which allows DP-bounded leakage during the computation.
 Typically, utilizing the computational DP model to leak intermediate steps implies sacrificing privacy or utility for an efficient runtime.
-However, in our case, we specifically choose to leak intermediate values that are typically accounted for in the central DP proof of DP-Lloyd (Blum et al., 2005; Su et al., 2016).
+However, in our case, we specifically choose to leak intermediate values already accounted for in the central DP proof of DP-Lloyd [9, 64].
 In other words, we get this speed-up with no additional cost to the privacy parameter $\epsilon$.
-Furthermore, rather than decrease the utility, we improve the utility of the clustering over state-of-the-art approaches (Su et al., 2016).
-To perform the aggregation over multiple parties, we design a modified secure aggregation protocol that keeps the aggregator oblivious to the client’s inputs and the centroids that are output at each iteration of the protocol.
-This allows us to add DP noise equivalent to the central model of DP.
-The combination of our improved DP algorithm, leaking the DP centroids, and our lightweight secure aggregation protocol allows us to reduce the computation time by up to four orders of magnitude compared to related work (Mohassel et al., 2020) (from minutes to a fraction of a second per iteration).
-In summary, our contributions include:
+Furthermore, rather than decrease the utility, we improve the utility of the clustering over state-of-the-art approaches [64].
+To perform the aggregation over multiple parties, we design a lightweight secure aggregation protocol that keeps the aggregator oblivious to the client’s inputs and the global centroids that are output at each iteration of the protocol.
+This allows the aggregator to add DP noise equivalent to the central model of DP.
+The combination of our improved DP algorithm, leaking the DP centroids, and our lightweight, secure aggregation protocol allows us to reduce the computation time by up to five orders of magnitude compared to related work [50] (from minutes to milliseconds per iteration).
+In summary, our contributions are four-fold:
   • We design the first DP protocol for horizontally federated private $k$-means.
-  • We improve the clustering utility over state-of-the-art DP $k$-means solutions by using cluster-size constraints and centroid-level aggregation.
-  • We design an efficient protocol using DP and a modified secure aggregation protocol to implement our protocol in the local model of trust.
-  • We prove our protocol is secure and preserves the end-to-end privacy in the computational model of DP and reduces the runtime by four orders of magnitude over the state-of-the-art secure federated approaches.
+  • We improve the clustering utility over state-of-the-art DP $k$-means solutions by developing a new DP algorithm with various improvements, such as enforcing a radius constraint on the centroids and using relative cluster updates.
+  • We design an efficient protocol using DP and a lightweight secure aggregation protocol to implement our protocol in the local trust model.
+  • We prove our protocol is secure and preserves the end-to-end privacy in the computational model of DP and reduces the runtime by five orders of magnitude over the state-of-the-art secure federated approaches.
 The remainder of the paper is organized as follows.
@@
 We then present our complete protocol for federated DP $k$-means (FastLloyd) and prove its privacy and utility in Section 5.
-In Section 6, we give an in-depth evaluation of our protocol in terms of utility, runtime, and communication size over various real-world and synthetic datasets.
-Finally, Section 7 describes extensions to our protocol, including a distance-constrained method and alternatives to the aggregation server.
-2. Background
-2.1. Notation
+Finally, in Section 6, we give an in-depth evaluation of our protocol in terms of utility, runtime, and communication size over various real-world and synthetic datasets.
+2 Background
+2.1 Notation
 Throughout this paper, we will denote objects that we intend to further slice/index with boldface notation (e.g.
-$\boldsymbol{\mu}$, while keeping atomic objects (whether scalars or vectors) as normal face (e.g.
+$\bm{\mu}$), while keeping atomic objects (whether scalars or vectors) as normal face (e.g.
 $\epsilon$).
-A summary of the notation used in this paper is provided in Table 1 and we will define the notation as it is used.
+A summary of the notation used in this paper is provided in Table 1, and we will define the notation as it is used.
 Symbol Description
-$\boldsymbol{p}$ Set of $M$ clients: $\{p_{1},p_{2},...,p_{M}\}$
+$\bm{p}$ Set of $M$ clients: $\{p_{1},p_{2},...,p_{M}\}$
 S Service provider
 $D_{i}$ Local dataset of client $p_{i}$: $\{x_{1},x_{2},...,x_{N_{i}}\}$
-$N_{i}$ Size of the dataset of client $p_{i}$
 $d$ Dimension of the data points
@@
 $\mathbf{O}$ Set of $k$ clusters: $\{O_{1},O_{2},...,O_{k}\}$
-$\boldsymbol{\mu}$ Centroids of clusters: $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
-$\boldsymbol{\gamma}$ Differentially-private noise
-$\eta$ Constraint parameter
-$\epsilon$ Privacy budget
-Table 1. Notation used in the paper.
-2.2. $k$-means Lloyd’s Algorithm
+$\bm{\mu}$ Centroids of clusters: $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
+$S_{j}$ Sum of data points in cluster $j$
+$C_{j}$ Count of data points in cluster $j$
+$R_{j}$ Relative sum of data points in cluster $j$
+$\bm{\gamma}$ Differentially-private noise
+$\eta$ Maximum radius bound
+$(\epsilon,\delta)$ Privacy budget
+$\sigma$ Noise multiplier
+Table 1: Notation used in the paper.
+2.2 $k$-Means Problem
 The $k$-means problem is a discrete optimization problem that aims to partition a set of $N$ $d$-dimensional observations into $k$ clusters, each represented by its mean or centroid.
-Given a dataset of observations $D=\{x_{1},x_{2},...,x_{N}\}$, where each observation is a $d$-dimensional real vector, the $k$-means problem is to find an assignment of data points to clusters, and a set of cluster centers, that minimizes the Within-cluster Sum of Squares (WCSS) objective:
-(1) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
-where $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$, $O_{\subseteq}D$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the centroids of the clusters, defined to be the (arithmetic) mean of points in $O_{j}$.
-Lloyd’s Algorithm
-The $k$-means problem is NP-hard in Euclidean space (Aloise et al., 2009).
-However, a simple heuristic known as Lloyd’s algorithm is commonly applied for practical applications (Lloyd, 1982).
+Given a dataset of observations $D=\{x_{1},x_{2},...,x_{N}\}$, where each observation is a $d$-dimensional real vector, the $k$-means problem is to find an assignment of data points to clusters, and a set of cluster centroids, that minimizes the Within-cluster Sum of Squares (WCSS) objective:
+$\underset{\mathbf{O},\bm{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}_{2}$ (1)
+where $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$, $O_{j}\subseteq D$ are the clusters, and $\bm{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the centroids of the clusters, defined to be the (arithmetic) mean of points in $O_{j}$.
+2.2.1 Lloyd’s Algorithm
+The $k$-means problem is NP-hard in Euclidean space [2].
+However, a simple heuristic known as Lloyd’s algorithm is commonly applied for practical applications [44].
 We detail the basic procedure of Lloyd’s algorithm:
-  1. (1)
+  1. 1.
     Initialization step: Randomly sample $k$ centroids $\{\mu_{1},\mu_{2},...,\mu_{k}\}$ (typically from the datapoints).
-  2. (2)
+  2. 2.
     Repeat until convergence (the assignments no longer change) or a predetermined number of iterations has been reached:
@@
       Assignment step: Assign each observation $x_{l}$ to the nearest centroid (using the Euclidean distance). This creates clusters $O_{j}$ for $j=1,2,...,k$. Formally, the assignment is:
-      $O_{j}^{(t)}=\big{\{}x_{l}:||x_{l}-\mu_{j}^{(t-1)}||^{2}\leq||x_{l}-\mu_{c}^{(t-1)}||^{2}\ \forall c\in\{1,\dots,k\}\}$
+      $O_{j}^{(t)}=\big{\{}x_{l}:||x_{l}-\mu_{j}^{(t-1)}||^{2}_{2}\leq||x_{l}-\mu_{c}^{(t-1)}||^{2}_{2}\ \forall c\in\{1,\dots,k\}\}$
     2. (b)
       Update step: Calculate the new centroids to be the mean of the observations in the cluster:
-      $\mu_{j}^{(t)}=\frac{\sum\limits_{x_{l}\in\mathbf{O}{j}^{(t)}}x_{l}}{\left|O_{j}^{(t)}\right|}$
-2.3. Constrained $k$-means
-The constrained $k$-means problem is a variant of the $k$-means problem where, in addition to minimizing the WCSS (Eqn 1), a certain set of constraints $\boldsymbol{\eta}$ must be satisfied.
-The problem was first introduced by Bennet et.
-al (Bennett et al., 2000), where they propose a method to enforce minimum cluster size constraints, $\boldsymbol{\eta}^{min}$.
-To solve the constrained $k$-means problem, Lloyd’s algorithm can be modified to incorporate the constraints in the assignment step of every iteration.
-However, the assignment step becomes nontrivial as it requires the solution to a constrained optimization problem.
-Bennet et.
-al (Bennett et al., 2000) provide an efficient solution to this constrained optimization problem by formulating it as a min-cost flow problem, which can be solved using off-the-shelf network flow solvers.
-An open-source implementation by Levy-Kramer (Levy-Kramer, 2018) implements the approach, using Google’s ORTools (Perron and Furnon, 2023) to solve the flow problem, and extends it to allow for maximum cluster size constraints, $\boldsymbol{\eta}^{max}$ (Eqn 2).
-(2) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}\quad\text{s.t.}\quad\eta_{j}^{min}\leq|O_{j}|\leq\eta_{j}^{max}\quad\forall j$
-Figure 1 shows an example reduction of an assignment instance to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
-[…]
-Figure 1. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
-Datapoints are represented as nodes and connected to cluster nodes through directed edges.
-The cost of each edge $(x_{l}\rightarrow O_{j})$ is defined to be the Euclidean distance between $x_{l}$ and the corresponding centroid $\mu_{j}$.
-The capacity of each edge is set to $1$, and each data node has a supply of $1$, ensuring that each data point is associated with exactly one cluster.
-On the other hand, each centroid node has a supply of $-\eta_{min}$ (equivalently, a demand of $\eta_{min}$).
-As every valid flow solution must satisfy this demand, a minimum of $\eta_{min}$ data points will flow into each cluster.
-To impose the maximum size constraint, clone nodes are created for each centroid, and each centroid is connected with its clone through an edge with zero cost and a capacity of $\eta_{max}$.
-This constrains the maximum flow into each cluster to be $\eta_{max}$.
-Lastly, a sink node is introduced to consume the remaining flow.
-A valid flow solution respects the constraints on cluster sizes, and by minimizing the cost, the total distance of points to their assigned centroids is minimized.
-Therefore, a min-cost flow solution to this model solves the constrained $k$-means problem for both minimum and maximum cluster constraints.
-2.4. Differential Privacy
-Differential privacy (DP) (Dwork et al., 2006) is an increasingly popular notion for protecting individuals’ privacy while allowing the computation of aggregate statistics.
+      $\mu_{j}^{(t)}=\frac{\sum\limits_{x_{l}\in O_{j}^{(t)}}x_{l}}{\left|O_{j}^{(t)}\right|}$
+2.3 Differential Privacy
+Differential privacy (DP) [17] is an increasingly popular notion to protect the privacy of individuals while allowing the computation of aggregate statistics.
 Differential privacy guarantees that an algorithm’s output is approximately the same, regardless of the participation of any single user.
@@
 Definition 2.1 (Differential Privacy).
-A randomized algorithm $M:\mathcal{D}\mapsto\mathbb{R}$ is $\epsilon$-DP, if for any pair of neighbouring datasets $D,D^{\prime}\in\mathcal{D}$, and for any $T\subseteq\mathbb{R}$ we have
-(3) $\Pr[M(D)\in T]\leq e^{\epsilon}\Pr[M(D^{\prime})\in T].$
-The privacy parameter $\epsilon$ defines how similar the outputs must be and determines the trade-off between privacy and utility.
+A randomized algorithm $M:\mathcal{D}\mapsto\mathbb{R}$ is $(\epsilon,\delta)$-DP, if for any pair of neighbouring datasets $D,D^{\prime}\in\mathcal{D}$, and for any $T\subseteq\mathbb{R}$ we have
+$\Pr[M(D)\in T]\leq e^{\epsilon}\Pr[M(D^{\prime})\in T]+\delta.$ (2)
+The privacy parameter $\epsilon$ defines how similar the outputs must be, and $\delta$ allows a small chance of failure in the definition.
 We use the unbounded neighbouring definition where datasets are neighbours if $|D\backslash D^{\prime}\cup D^{\prime}\backslash D|=1$.
-That is, we allow the addition or removal of a single data point.
-We note that arbitrary computations can be carried out on the output of a DP mechanism without affecting privacy (the postprocessing lemma (Dwork and Roth, 2014)).
+That is, we allow for the addition or removal of a single data point.
+We note that arbitrary computations can be carried out on the output of a DP mechanism without affecting privacy (the post-processing lemma [18]).
 Finally, DP is composed naturally with multiple runs of a mechanism.
-If we apply a differentially private mechanism(s) sequentially, the privacy parameters are composed through summation or more advanced methods (Dwork and Roth, 2014).
-If a mechanism is applied multiple times over disjoint subsets of the dataset, then the total privacy leakage is the maximum privacy parameter over each subset (parallel composition (Dwork and Roth, 2014)).
+If we apply a differentially private mechanism(s) sequentially, the privacy parameters are composed through summation or more advanced methods [18].
+If a mechanism is applied multiple times over disjoint subsets of the dataset, then the total privacy leakage is the maximum privacy parameter over each subset (parallel composition [18]).
 Definition 2.2 (Sensitivity).
 Let $f:\mathcal{D}\mapsto\mathbb{R}^{k}$. If $\ \mathbb{D}$ is a distance metric between elements of $\ \mathbb{R}^{k}$ then the $\mathbb{D}$-sensitivity of $f$ is
-(4) $\Delta^{(f)}_{d}=\max_{(D,D^{\prime})}\mathbb{D}(f(D),f(D^{\prime})),$
+$\Delta^{(f)}=\max_{(D,D^{\prime})}\mathbb{D}(f(D),f(D^{\prime})),$ (3)
 where $(D,D^{\prime})$ are pairs of neighbouring datasets.
-Common examples of the distance metric are the $\ell_{1}$-norm and $\ell_{2}$-norm.
-With the notion of sensitivity, we can now define the Laplace mechanism, which uses sensitivity based on the $\ell_{1}$-norm, which we denote $\Delta^{(f)}_{1}$ for simplicity.
-Definition 2.3 (Laplace Mechanism).
-Let $f:\mathcal{D}\mapsto\mathbb{R}^{k}$. The Laplace mechanism is defined as
-(5) $M(D)=f(D)+\boldsymbol{\gamma},$
-where
-$\boldsymbol{\gamma}\in\mathbb{R}^{k}\quad\text{s.t.}\quad\gamma_{i}\overset{\mathrm{iid}}{\sim}Lap(\Delta^{(f)}_{1}/\epsilon)$
-i.e. $\boldsymbol{\gamma}$ represents $k$ i.i.d draws from the Laplace distribution with scale parameter $\Delta^{(f)}_{1}/\epsilon$.
-It is well known that the Laplace mechanism satisfies $\epsilon$-DP (Dwork and Roth, 2014).
-3. Problem Statement
-We consider $M$ clients denoted by $\boldsymbol{p}=\{p_{1},p_{2},...,p_{M}\}$, with each party $p_{i}$ owning a private dataset $D_{i}=\{x_{1},x_{2},...,x_{N_{i}}\}$, where $x_{l}\in\mathbb{R}^{d}$, $d$ denotes the dimensionality of the dataset and $N_{i}$ is the size of the dataset of party $i$.
-For simplicity, we assume that datasets held by each client are of similar sizes.
-The objective is to compute a collaborative $k$-means clustering $(\mathbf{O},\boldsymbol{\mu})$ over $\boldsymbol{D}$ that minimizes the WCSS objective (1); where $\boldsymbol{D}=\bigcup_{i=1}^{i=M}D_{i}$ is the union of the datasets held by the clients, $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the respective centroids of the clusters.
+We will focus on the $\ell_{2}$ norm in this work as we use the Gaussian Mechanism. To analyze the Gaussian Mechanism, we will use Gaussian Differential Privacy (GDP) [15].
+Definition 2.3 (GDP [15]).
+A mechanism $M$ is said to satisfy $\theta$-Gaussian Differential Privacy ($\theta$-GDP) if it is $G_{\theta}$-DP. That is,
+$T\big{(}M(D),M(D^{\prime})\big{)}\geq G_{\theta}$
+for all neighboring datasets $D$ and $D^{\prime}$, where $T$ is a trade-off function and $G_{\theta}=T\big{(}\mathcal{N}(0,1),\mathcal{N}(\theta,1)\big{)}$ (see Dong et al. [15] for more details).
+Naturally, the Gaussian Mechanism satisfies GDP.
+Theorem 2.1.
+(Gaussian Mechanism GDP [15]) Define the Gaussian mechanism that operates on a statistic $f$ as $M(D)=f(D)+\gamma$, where $\gamma\sim\mathcal{N}(0,(\Delta^{(f)})^{2}/\theta^{2})$.
+Then, $M$ is $\theta$-GDP.
+Similar to DP, GDP composes over multiple adaptive uses of a mechanism.
+Theorem 2.2 (GDP Composition [15]).
+The $n$-fold composition of $\theta_{i}$-GDP mechanisms is $\sqrt{\theta_{1}^{2}+\cdots+\theta_{n}^{2}}$-GDP.
+Finally, it is possible to convert a GDP guarantee to DP and vice versa:
+Theorem 2.3 (GDP to DP [15, 6]).
+A mechanism is $\theta$-GDP if and only if it is $\big{(}\epsilon,\delta(\epsilon)\big{)}$-DP for all $\epsilon\geq 0$, where
+$\delta(\epsilon)=\Phi\Big{(}-\frac{\varepsilon}{\theta}+\frac{\theta}{2}\Big{)}-e^{\epsilon}\Phi\Big{(}-\frac{\varepsilon}{\theta}-\frac{\theta}{2}\Big{)}.$
+In practice, we use the algorithm derived by Balle and Wang to solve this function for $\theta$ [6, Algorithm 1].
+3 Problem Statement
+We consider $M$ clients denoted by $\bm{p}=\{p_{1},p_{2},...,p_{M}\}$, with each party $p_{i}$ owning a private dataset $D_{i}=\{x_{1},x_{2},...,x_{N_{i}}\}$, where $x_{l}\in\mathbb{R}^{d}$, $d$ denotes the dimensionality of the dataset and $N_{i}$ is the size of the dataset of party $i$.
+The total dataset that we compute on, can be very large, even if there are only a small number of clients.
+The objective is to compute a collaborative $k$-means clustering $(\mathbf{O},\bm{\mu})$ over $\bm{D}$ that minimizes the WCSS objective (1); where $\bm{D}=\bigcup_{i=1}^{i=M}D_{i}$ is the union of the datasets held by the clients, $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$ are the clusters, and $\bm{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the respective centroids of the clusters.
 We aim to optimize this objective while formally proving the privacy of the output and intermediate computations.
 Specifically, the protocol should be secure in the computational model of differential privacy introduced by Mirnov et al.
-(Mironov et al., 2009) and extended to the multi-party setting by Humphries et al.
-(Humphries et al., 2022).
-Definition 3.1 (IND-CDP-MPC (Humphries et al., 2022; Mironov et al., 2009)).
-A multi-party protocol $\Pi$ for computing function $f$ satisfies $\epsilon(\lambda)$- indistinguishable computationally differential privacy (IND-CDP-MPC) if for every probabilistic polynomial time (in $\lambda$) adversary $\mathcal{A}$ with input $D_{A}$, and for neighbouring datasets $D,D^{\prime}$ belonging to the honest parties (i.e.
+[48] and extended to the multi-party setting by Humphries et al.
+[33].
+We further extend this model to the approximate DP setting by incorporating the failure probability $\delta$.
+Definition 3.1 (IND-CDP-MPC [33, 48]).
+A multi-party protocol $\Pi$ for computing function $f$ satisfies $(\epsilon(\lambda),\delta)$- indistinguishable computationally differential privacy (IND-CDP-MPC) if for every probabilistic polynomial time (in $\lambda$) adversary $\mathcal{A}$ with input $D_{A}$, and for neighbouring datasets $D,D^{\prime}$ belonging to the honest parties (i.e.
 $D,D^{\prime}=\cup_{i\setminus A}D_{i}$),
 $\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D))=1]$
-$\leq$ $\exp(\epsilon)\cdot\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D^{\prime}))=1]+negl(\lambda).$
+$\leq$ $\exp(\epsilon)\cdot\Pr[\mathcal{A}(\textsc{view}_{A}^{\Pi}(D_{A},D^{\prime}))=1]+negl(\lambda)+\delta.$
 Likewise, for every other party’s view of neighbours $(D,D^{\prime})$.
 Intuitively, after observing the output and intermediate computations, the amount of information that any party can learn about any other party’s local dataset should be bounded by differential privacy.
+Compared to the local and shuffle models of DP, the IND-CDP-MPC model provides a significantly better privacy vs.
+utility trade-off at the level of the central model of DP.
+We show that removing the need for a trusted aggregator (needed in the central DP model) does not require the computational resources of a large-scale secure computation protocol.
+We achieve this by processing data in encrypted form (secure computation), accelerated by revealing differentially private intermediate values.
 Finally, we make the following assumptions when designing our protocol:
   • The existence of an honest-but-curious service provider S to assist with multiparty computations.
-This party should be oblivious to all inputs and results, i.e., the service provider learns nothing about the client’s input or output during the execution of the protocol.
-We discuss alternatives to this assumption in Section 7.
-  • Clients share a common secret before the start of the protocol; this can be used as a seed to a common Pseudo-Random Number Generator (PRNG) and is generally possible through public-key schemes or communication over a channel secured by TLS.
-4. Related Work
-4.1. Secure Exact $k$-means
-We first discuss related work using secure computation to compute $k$-means that does not preserve the privacy of the output.
+This party should be oblivious to all inputs and results, i.e., the service provider learns nothing about the client’s input or output during the execution of the protocol (a weaker assumption than the local and shuffle models).
+We also discuss alternatives to this assumption in Appendix C.
+  • Clients share a common secret used as a PRNG seed and do not collude with the server.
+This secret can be established through standard public-key schemes (e.g., Bonawitz et al.
+[10] uses Diffie-Hellman key exchange [14]) or TLS-secured channels.
+4 Related Work
+4.1 Secure Exact $k$-means
+We first discuss secure computation based approaches to $k$-means that do not preserve the privacy of the output.
 This literature shows a trade-off between privacy and efficiency.
-Approaches that prioritize efficiency often compromise privacy by allowing the leakage of intermediate computations such as sums, counts, or centroids, which can be exploited to infer sensitive data (Gheid and Challal, 2016; Jiang et al., 2020).
+Approaches that prioritize efficiency often compromise privacy by allowing the leakage of intermediate computations such as sums, counts, or centroids, which can be exploited to infer sensitive data [29, 38].
 For instance, Gheid et al.
-(Gheid and Challal, 2016) rely on computing secure sums for aggregating sums and counts across clients.
-While this is very efficient in practice, it reveals aggregate intermediate values to all clients, which is not considered secure (Hegde et al., [n.
-d.]).
+[29] rely on computing secure sums for aggregating sums and counts across clients.
+While this is very efficient in practice, it reveals aggregate intermediate values to all clients, which is not considered secure [31].
+As we discuss in Appendix C, naively fixing this leakage would require a significant degradation in utility or performance.
 Another approach by Jiang et al.
-(Jiang et al., 2020) employs homomorphic encryption and garbled circuits to conceal sums and counts.
+[38] employs homomorphic encryption and garbled circuits to conceal sums and counts.
 However, it still reveals the intermediate centroids to all clients in the clear.
 Many other works in this category are explained in the SoK of Hedge et al.
-(Hegde et al., [n.
-d.]).
-Conversely, secure approaches that do not leak intermediate computations employ heavy cryptographic techniques, like fully homomorphic encryption, leading to significant performance degradation (Bunn and Ostrovsky, 2007; Rao et al., 2015; Jäschke and Armknecht, 2018; Kim and Chang, 2018).
-Bunn and Ostrovsky (Bunn and Ostrovsky, 2007) utilize arithmetic secret sharing alongside homomorphic encryption, with high-performance overheads.
+[31].
+Conversely, secure approaches that do not leak intermediate computations employ heavy cryptographic techniques, like fully homomorphic encryption, leading to significant performance degradation [11, 59, 36, 40].
+Bunn and Ostrovsky [11] utilize arithmetic secret sharing alongside homomorphic encryption, with high-performance overheads.
 Homomorphic encryption has also been used by Rao et al.
-(Rao et al., 2015), Jäschke et al.
-(Jäschke and Armknecht, 2018), and Kim et al.
-(Kim and Chang, 2018), and consistently results in prohibitively high computation time due to the computation demands of the encryption.
+[59], Jäschke et al.
+[36], and Kim et al.
+[40], and consistently results in prohibitively high computation time due to the computation demands of the encryption.
 The work of Mohassel et al.
-(Mohassel et al., 2020) is a notable exception, offering a scheme that is significantly faster than previous state-of-the-art methods.
+[50] is a notable exception, offering a scheme that is significantly faster than previous state-of-the-art methods.
 By using optimized $2$-party computation primitives, batched oblivious transfers, and garbled circuits, they achieve a speedup of five orders of magnitude over Jäschke et al.
-(Jäschke and Armknecht, 2018).
+[36].
 However, the runtime is still on the order of minutes; and the communication size is on the order of gigabytes.
 In contrast, our approach is as efficient as the protocols that leak intermediate computations, while protecting the privacy of the output and intermediate computations with DP.
-4.2. Differentially Private $k$-means
+4.2 Differentially Private $k$-means
 DP $k$-means algorithms aim to address output privacy by introducing noise during clustering, hiding individual data contributions, and ensuring a formal privacy guarantee.
 Research in DP $k$-means is divided into central and local models.
-Central DP
+4.2.1 Central DP
 The central model operates with a trusted curator who collects data for analysis, hiding individuals’ information in the output, but not from the curator itself.
-Among the central approaches, DP-Lloyd introduces noise during each centroid update, by adding Laplace noise to both the sum (numerator) and the counts (denominator) when computing the arithmetic mean of points within a cluster (Blum et al., 2005).
-DP-Lloyd was later implemented in the PinQ framework (McSherry, 2009) with a fixed number of iterations.
-Dwork (Dwork, 2011) extended this framework to allow for an arbitrary number of iterations, allocating an exponentially decreasing privacy budget to each iteration, although it was noted that utility degrades beyond a certain number of iterations due to increasing noise levels.
+Among the central approaches, DP-Lloyd introduces noise during each centroid update, by adding Laplace noise to both the sum (numerator) and the counts (denominator) when computing the arithmetic mean of points within a cluster [9].
+DP-Lloyd was later implemented in the PinQ framework [47] with a fixed number of iterations.
+Dwork [16] extended this framework to allow for an arbitrary number of iterations, allocating an exponentially decreasing privacy budget to each iteration, although it was noted that utility degrades beyond a certain number of iterations due to increasing noise levels.
 The state-of-the-art in DP-Lloyd was achieved by Su et al.
-(Su et al., 2016), who included an error analysis to determine the optimal number of iterations and introduced a “sphere-packing” centroid initialization method that significantly improves clustering quality.
-The Sample and Aggregate Framework (SaF) partitions the dataset into multiple subsets, upon which the non-private Lloyd’s algorithm is executed independently (Nissim et al., 2007).
+[64], who included an error analysis to determine the optimal number of iterations, optimized the splitting of privacy budget between sum and count queries, and introduced a “sphere-packing" centroid initialization method that significantly improves clustering quality.
+The Sample and Aggregate Framework (SaF) partitions the dataset into multiple subsets, upon which the non-private Lloyd’s algorithm is executed independently [53].
 The resulting centroids from each subset are then aggregated using a standard DP mechanism.
-SaF was later implemented within the GUPT system (Mohan et al., 2012).
+SaF was later implemented within the GUPT system [49].
 However, experiments conducted by Su et al.
-(Su et al., 2016) suggest that DP-Lloyd consistently outperforms SaF across a variety of synthetic and real datasets.
+[64] suggest that DP-Lloyd consistently outperforms SaF across various synthetic and real datasets.
 Synopsis-based methods take a different approach by first generating a synopsis of the dataset using a differentially private algorithm, then applying the $k$-means clustering algorithm on this synopsis.
 Qardaji et al.
-(Qardaji et al., 2012) proposed a synopsis technique for 2D datasets, which was extended and optimized by Su et al.
-(Su et al., 2016) to higher dimensions.
-However, this approach only outperforms DP-Lloyd in datasets with less than three dimensions (Su et al., 2016).
-Another, more theoretical, line of work focuses on minimizing approximation error bounds, without implementation or empirical evaluation (Feldman et al., 2009, 2017; Balcan et al., 2017; Nissim and Stemmer, 2018; Stemmer and Kaplan, 2018; Ghazi et al., 2020).
+[57] proposed a synopsis technique for 2D datasets, which was extended and optimized by Su et al.
+[64] to higher dimensions.
+However, this approach only outperforms DP-Lloyd in datasets with less than three dimensions [64].
+Another, more theoretical, line of work focuses on minimizing approximation error bounds, without implementation or empirical evaluation [20, 21, 3, 54, 63, 27].
 The work of Ghazi et al.
-(Ghazi et al., 2020) concludes this long line of work by achieving the same approximation factor as the best non-private algorithms.
+[27] concludes this long line of work by achieving the same approximation factor as the best non-private algorithms.
 However, these methods suffer from superlinear running times, making them impractical for large datasets, where privacy is most important.
-Local DP
+4.2.2 Local DP
 In contrast to central DP, the local model operates under the premise that no trusted curator is available.
-This necessitates that individuals randomize their data locally before it is aggregated, which reduces utility by introducing a significantly higher level of noise (on the order of the number of clients).
-Most research in this area, including works by Nissim and Stemmer (Nissim and Stemmer, 2018) and Stemmer (Stemmer, 2021), focus on theoretical error bounds rather than practical applications.
+This necessitates that individuals randomize their data locally before it is aggregated, which reduces utility by introducing a significantly higher level of noise (proportional to the number of data points).
+Several works have been proposed in this area [54, 62, 13], most of which focus on the theoretical error bounds rather than practical applications.
+Another line of work studies private aggregation (a building block for federated DP $k$-means) in local and shuffle DP models [28, 4].
+However, these works cannot overcome the utility gap stemming from the noise being added to each data point.
 Finally, Li et al.
-(Li et al., 2023) attempt to bridge the gap between these central and local models in the vertically federated setting.
+[43] attempt to bridge the gap between these central and local models in the vertically federated setting.
 Their approach relies on an untrusted curator aggregating noisy local centers and membership encodings from the clients to generate a private synopsis.
 However, this approach is specifically tailored for the vertical setting and cannot be applied to our configuration.
-Our work yields utility that is even better than state-of-the-art work in the central model (Su et al., 2016) while obtaining the trust assumptions of the local model and operating in the federated setting.
-5. Distributed DP $k$-means
-First, we overview our modified DP algorithm, using constrained $k$-means.
+Our work yields utility that is even better than state-of-the-art work in the central model [64] while obtaining the trust assumptions of the local model and operating in the federated setting.
+5 Distributed DP $k$-means
+First, we overview our modified DP algorithm using the maximum radius constraint and relative updates.
 Then, we describe FastLloyd, including the initialization, assignment, and update steps.
-Finally, we provide an error analysis and proof of privacy.
-5.1. Constrained DP $k$-means
-We observe that neither federated exact $k$-means nor DP $k$-means algorithms in the literature take advantage of constrained clustering approaches from the non-private literature (Bennett et al., 2000; Levy-Kramer, 2018).
-In this section, we describe how incorporating constrained clustering into the assignment step significantly improves utility; a contribution that is independently applicable to the central model of DP.
-In the DP $k$-means literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), the typical way to compute a differentially-private mean is to perform two differentially-private queries: one for the sum and one for the count.
-The sum is then divided by the count as a postprocessing step.
-We call this the sum-count method.
-If the datapoints $x_{l}$ are contained in $[-B,B]$, then the sensitivity of the sum is $B$, and the sensitivity of the count is $1$.
-Adding noise at these sensitivities yields reasonable utility, since the sum has an expected magnitude much larger than $B$, so the noise does not dominate the value, and upon division, the noise is further reduced.
-The sum-count method is typically more privacy-efficient than releasing the mean directly.
-If we release the mean directly, the sensitivity needs to account for the existence of clusters with a single data point, leading to a sensitivity of $B$ for the mean.
-Since the magnitude of the mean itself is at most $B$, a sensitivity of $B$ produces a relatively large amount of noise, leading to unusable utility.
-We show that by using constrained clustering, we can tighten the sensitivity bound and release the mean directly.
-This avoids dividing the privacy budget between the sum and count queries and gives a much higher utility than the sum-count method as we show in Section 6.2.
-By restricting the minimum cluster size to $\eta_{min}$, we provide a tunable bound on the sensitivity of the mean.
-Since all clusters have at least $\eta_{min}$ data points, the denominator of the centroid is now bounded below by $\eta_{min}$.
-Therefore, a single data point cannot affect any centroid by more than $\frac{B}{\eta_{min}}$.
-(6) $\Delta^{\mu}=\frac{B}{\eta_{min}}$
-The centroid-method for computing the DP mean is:
-(7) $\bar{\mu_{j}}=\frac{\sum_{x_{l}\in O_{j}}x_{l}}{|O_{j}|}+\gamma^{\mu}\quad\text{where}\quad\gamma^{\mu}\sim Lap\left(\frac{\Delta^{\mu}}{\epsilon^{\mu}}\right)$
-We further claim that, by employing constrained clustering, the sum-count method can also be improved.
-Intuitively, having a minimum cluster size prevents the formation of very small clusters, where the differentially-private noise would dominate the underlying sum and count.
-In Section 6.2, we show that the sum-count method is improved by using constrained clustering, but not as much as the centroid method.
-5.2. Overview of FastLloyd
+Finally, we provide our security theorem, and discuss how we choose the various parameters.
+5.1 Radius Constrained DP $k$-means
+In this section, we describe how we improve the clustering utility over the work of Su et al.
+[64], a contribution that is independently applicable to the central model of DP.
+We observe that despite being common in the non-private literature [8, 42], neither federated exact $k$-means nor DP $k$-means algorithms in the literature apply any constraints to the $k$-means objective.
+In this work, we focus on a type of constraint not studied in the non-private literature: bounding the radius of each cluster.
+This constraint is particularly useful in the context of DP $k$-means, as it allows a tighter sensitivity bound when computing cluster updates relative to the previous centroid.
+In the DP $k$-means literature [9, 47, 16, 64], a typical private clustering algorithm is to run Lloyd’s algorithm, with DP noise added to the update step.
+The DP mean is computed by two DP queries: one for the sum and one for the count.
+The sum is then divided by the count as a post-processing step.
+The challenge in these protocols is that the sensitivity of the sum is bounded by the domain size.
+If the datapoints $x_{l}$ are contained in $[-B,B]^{d}$, then the $l_{2}$ sensitivity of the sum is $B\sqrt{d}$.
+This is because, with no other constraints, there exists a worst-case data point that could add $B$ to the sum in all dimensions, regardless of the cluster it is assigned to.
+In this work, we modify the $k$-means objective such that in addition to minimizing the WCSS (Eqn 1), a constraint on the maximum radius of each cluster must also be satisfied.
+Specifically, no data point in a cluster can be more than $\eta$ away from the cluster’s centroid.
+Thus, the objective becomes to minimize (Eqn 1) s.t.:
+$||x_{l}-\mu_{j}||_{2}\leq\eta\quad\forall x_{l}\in O_{j},\forall j$ (4)
+To enforce this constraint, we modify the assignment step to not assign a data point to a cluster if it is more than $\eta$ away from its nearest centroid.
+Any unassigned data points are discarded and, therefore, do not contribute to any sum or count query.
+Intuitively, bounding the maximum radius of each cluster limits how much a cluster can move by adding or removing a data point since a worst-case data point must be close to a cluster’s centroid to be factored into the mean.
+However, simply bounding the radius of the cluster does not tighten the sensitivity of the sum query.
+This is because the sensitivity analysis must consider the worst-case cluster.
+If a cluster is within $\eta$ of the domain boundary, then the $\ell_{2}$ sensitivity of the sum query is still $B\sqrt{d}$.
+To realize the reduction in sensitivity from bounding the radius, we must modify the update step as well as the assignment step.
+Instead of perturbing the sum of the data points themselves, we perturb the difference between the data points and their assigned cluster centroid.
+Specifically:
+$R_{j}^{(t)}=\sum\limits_{x_{l}\in O_{j}^{(t)}}x_{l}-\tilde{\mu}_{j}^{(t-1)}$ (5)
+Essentially, we are computing the “updates” to the centroids, rather than the centroids themselves.
+This simple change has a significant impact on the sensitivity of the sum query.
+The sensitivity of the relative sum is now bounded by the maximum radius parameter $\eta$:
+Theorem 5.1.
+If the constraint (4) is satisfied, then:
+$\Delta^{R}=\max\limits_{D,D^{\prime}\in\mathcal{D}}||R_{j}^{(t)}(D)-R_{j}^{(t)}(D^{\prime})||_{2}\leq\eta$ (6)
+for all clusters $j\in[k]$, and iterations $t\in[T]$.
+We prove this in Appendix A.1. The intuition is that a worst-case data point only contributes its distance to the centroid (which is bounded by $\eta$), rather than its distance to the origin.
+Our maximum distance constraint naturally bounds the $l_{2}$ sensitivity as it bounds the Euclidean distance, and thus, Gaussian noise is a natural choice that also scales more efficiently to higher dimensions.
+It also allows for a simple and tight privacy analysis using Gaussian-DP [15].
+We use the analysis of Balle and Wang [6, Algorithm 1] to obtain the noise multiplier $\sigma$.
+We divide the privacy budget between the relative sum and the count following our analysis in Section 5.4 as:
+$\sigma^{R}=\frac{\sigma\sqrt{1+\sqrt{4d}}}{\sqrt[4]{4d}}\quad\sigma^{C}=\sigma\sqrt{1+\sqrt{4d}}$ (7)
+We then add noise as follows:
+$\tilde{R}_{j}^{(t)}=R_{j}^{t}+\gamma^{R}\quad\text{where}\quad\gamma^{R}\sim\mathcal{N}(0,(\sigma^{R})^{2}\eta^{2}T)$ (8)
+$\tilde{C}_{j}^{(t)}=C_{j}^{t}+\gamma^{C}\quad\text{where}\quad\gamma^{C}\sim\mathcal{N}(0,(\sigma^{C})^{2}T).$ (9)
+We prove in Section 5.3 that adding noise in this way satisfies DP. After adding noise, we compute the new centroid as:
+$\tilde{\mu}_{j}^{(t)}=\frac{\tilde{R_{j}}^{(t)}}{\tilde{C_{j}}^{(t)}}+\tilde{\mu}_{j}^{(t-1)}.$ (10)
+Without DP noise, this yields an equivalent cluster update.
+$\mu_{j}^{(t)}$ $=$ $\frac{S_{j}^{(t)}-C_{j}^{(t)}\mu_{j}^{(t-1)}}{C_{j}^{(t)}}+\mu_{j}^{(t-1)}=\frac{S_{j}^{(t)}}{C_{j}^{(t)}}$ (11)
+where $S_{j}^{(t)}=\sum_{x_{l}\in O_{j}^{(t)}}x_{l}$ is the sum of the data points. With the DP noise, we get an additional noise (error) term in the sum compared to noising the sum directly:
+$\tilde{S}_{j}^{(t)}$ $=$ $\tilde{R}_{j}^{(t)}+\tilde{C}_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}$ (12)
+$=$ $S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}+\gamma^{R}+(C_{j}^{(t)}+\gamma^{C})\tilde{\mu}_{j}^{(t-1)}$ (13)
+$=$ $S_{j}^{(t)}+\gamma^{R}+\gamma^{C}\tilde{\mu}_{j}^{(t-1)}.$ (14)
+However, as we show in Section 6.2, this additional error is compensated for by the increase in utility from reducing the sensitivity from $B\sqrt{d}$ to $\eta$.
+The maximum radius further decreases error by reducing the effect of outliers, as data points far away from any centroid will not (and, in some cases, should not) be assigned to any cluster.
+We also use the maximum radius constraint as an additional post-processing constraint, which we call Radius Clipping.
+Namely, if a noisy centroid is computed to be more than $\eta$ away from the previous centroid, we truncate it to be $\eta$ away.
+5.2 Overview of FastLloyd
 In this section, we describe our protocol for federated DP $k$-means, FastLloyd.
+The focus of FastLloyd is to create the most efficient and accurate protocol possible under the threat model defined in Section 3.
+Specifically, we refrain from adding additional noise or computations that would be needed to handle client failures or collusion between clients and the server.
+We leave it to future work to adapt our approach to use more resilient aggregation protocols [10, 39, 61, 22].
 Algorithm 1 overviews the protocol from the perspective of a single client.
-Following the structure of Lloyd’s algorithm, the following sections describe how we design each of its main steps: initialization, assignment, and update.
-5.2.1. Initialization: Sphere Packing
+Following the outline of Lloyd’s algorithm, the following sections describe how we design each of its main steps: initialization, assignment, and update.
+5.2.1 Initialization: Sphere Packing
 We modify the initialization so that it can be carried out in a federated manner.
 We employ the sphere packing initialization approach of Su et al.
-(Su et al., 2016) as it was shown to outperform random initialization.
+[64] as it was shown to outperform random initialization.
 The sphere packing approach is data-independent and thus does not use any privacy budget.
 The process can be outlined as follows:
-  1. (1)
+  1. 1.
     Initialize a radius parameter $a$.
-  2. (2)
+  2. 2.
     For $j\in\{1,2,...,k\}$, generate a point $\mu_{j}$ such that it is at least of distance $a$ away from the domain boundaries and at least of distance $2a$ away from any previously chosen centroid.
 If a randomly generated point does not meet this condition, generate another one.
-  3. (3)
-    If, after repeated attempts, it is not possible to find such a point, decrease the radius $a$ and repeat the process.
+  3. 3.
+    If, after $100$ repeated attempts, it is not possible to find such a point, decrease the radius $a$ and repeat the process.
 The radius $a$ is determined via a binary search to find the maximum $a$ that allows for the generation of $k$ centroids.
 In Algorithm 1, each client independently calls the Initialization function (Line 3) with the same random seed, which results in each client starting with the same centroids.
-5.2.2. Assignment: Constrained $k$-means
-To set the minimum cluster ($\eta_{min}$) and maximum cluster ($\eta_{max}$) sizes, we apply a heuristic of setting them as a ratio of the average cluster as follows:
-(8) $\eta_{min}=\frac{N}{\alpha_{min}k}\quad\text{and}\quad\eta_{max}=\frac{\alpha_{max}N}{k}$
-where $\alpha_{min}$ and $\alpha_{max}$ are hyperparameters that control the minimum and maximum cluster sizes.
-In Appendix B, we show that $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ provide a good balance between utility and feasibility.
-We use the open-source implementation from Levy-Kramer (Levy-Kramer, 2018) to enforce the constraints in the assignment step.
-In Line 6 of Algorithm 1, the client computes their assignments locally using the min-cost flow solver.
+5.2.2 Assignment: Radius Constrained $k$-means
+In Line 7, each client locally computes the closest cluster to each of their data points.
+If the data point is within $\eta$ of the cluster’s centroid, it is assigned to that cluster (Line 9).
+Any points further than $\eta$ from the cluster’s centroid are not assigned to any cluster.
+We discuss how we set $\eta$ in Section 5.5.
 Algorithm 1 FastLloyd from $p_{i}$’s perspective
 1:Inputs: Local Dataset $D_{i}$.
-2:Output: Cluster Centres $\boldsymbol{\mu}^{(T)}$
-3:$\boldsymbol{\mu}^{(0)}=$ Initialization($seed$)
+2:Output: Cluster Centres $\tilde{\bm{\mu}}^{(T)}$
+[…]
 4:for $t\in\{1:T\}$ do
 5:     // Assignment Step
-6:     Compute $\mathbf{O}_{i}^{(t)}=$ MinCostFlowSolve($D_{i},\boldsymbol{\mu}^{(t-1)}$)
-7:     // Local Update
-8:     for $j\in\{1,\dots,k\}$ do
-9:         Compute $\bar{\mu}_{ij}^{(t)}=\frac{\sum_{x_{l}\in O_{ij}^{(t)}}x_{l}}{|O_{ij}^{(t)}|}$
-10:     // Global Update
-11:     $\widehat{\boldsymbol{\mu}}=$ GlobalMSA($\mathbf{\bar{\mu}}_{i}^{(t)}/M$, $\Delta$, $\epsilon^{\mu},seed,t$)
-12:     // Post Process Result
-13:     $\boldsymbol{\mu}^{(t)}=$ Fold($\widehat{\boldsymbol{\mu}}$)
-5.2.3. Local Update
-The client has already locally assigned each of their data points to a cluster, following the constraints, in the previous step (Line 6).
-Next, they compute the sum divided by the count of each cluster, using their local dataset.
-We call this the local update, and it follows the classic Lloyd’s algorithm exactly.
-The output of this step is a matrix (of dimensions $(k,d)$) containing the average of each dimension for each cluster on client $p_{i}$’s data.
-5.2.4. Global Update: Masked Secure Aggregation
-We modify the update step to release a perturbed version of the cluster centres to each client under the CDP security model.
-By publishing the previous noisy centroids, the assignment and local centroid computation can be performed locally by each client instead of using a secure computation protocol.
-In the global update step (Line 11), we privately aggregate (average) the local centroids of all clients, revealing the noisy global centroids for the next round.
-The clients scale their local centroids down by $M$, before calling the global update so that the sum of the local updates results in the average of the local centroids.
+6:     for $x_{l}\in D_{i}$ do
+[…]
+8:         if $||x_{l}-\tilde{\mu}_{j^{\prime}}^{(t-1)}||_{2}<\eta$ then
+9:              $O_{ij^{\prime}}^{(t)}\leftarrow x_{l}$
+10:     // Local Update
+11:     for $j\in\{1,\dots,k\}$ do
+12:         Compute $\bar{R}_{ij}^{(t)}=\sum\limits_{x_{l}\in O_{ij}^{(t)}}x_{l}-\tilde{\mu}_{j}^{(t-1)}$
+13:         Compute $\bar{C}_{ij}^{(t)}=|O_{ij}^{(t)}|$
+14:     // Global Update
+15:     $\widehat{\bm{U}}=$ GlobalMSA($\bar{R}_{ij}^{(t)}$, $\sigma^{R}$, seed, $t$)
+16:     $\widehat{\bm{C}}=$ GlobalMSA($\bar{C}_{ij}^{(t)}$, $\sigma^{C}$, seed, $t$)
+17:     // Post Process Result
+18:     for $j\in\{1,\dots,k\}$ do
+19:         $\widehat{\mu}_{j}=\frac{\widehat{\bm{U}}}{\widehat{\bm{C}}}+\tilde{\mu}_{j}^{(t-1)}$
+20:     $\tilde{\bm{\mu}}^{(t)}=$ Fold($\widehat{\bm{\mu}}$)
+5.2.3 Local Update
+The client has already locally assigned each of their data points to a cluster, following the constraints in the previous step (Line 9).
+Next, they compute the relative sum and the count for each cluster using their local dataset.
+We call this the local update step.
+The output of the relative sum (Line 12) and the count (Line 13) are two matrices of dimensions $(k\times d)$ and $(k\times 1)$ respectively.
+5.2.4 Global Update: Masked Secure Aggregation
+We modify the update step to release a perturbed version of relative sums and the counts to each client under the CDP security model.
+By publishing the previous noisy centroids, the assignment and local update computation can be performed locally by each client instead of using a secure computation protocol.
+In the global update step (Lines 15 and 16), we privately aggregate (sum) the local updates of all clients, perturb the result, and reveal the noisy global relative sums and counts for the next round.
 We describe our aggregation protocol in Figure 1.
 We call this protocol Masked Secure Aggregation (MSA).
-In MSA, similar to regular secure aggregation protocols (Bonawitz et al., 2017), a service provider S securely aggregates values from $M$ clients, all while oblivious to every client’s contributions.
+In MSA, similar to regular secure aggregation protocols [10], a service provider S securely aggregates values from $M$ clients, all while being oblivious to every client’s contributions.
 However, in MSA, the server is also oblivious to the result of the aggregation operation; the server only acts as an aggregator (who also adds DP noise) and is oblivious to both the input and output.
 Each client $p_{i}$ possesses a private value (a matrix $v_{i}\in\mathbf{R}^{k\times d}$, where $\mathbf{R}$ could be $\mathbb{Z}_{2^{64}}$).
-These clients aim to collectively compute the element-wise sum of their private matrices: $\boldsymbol{v}=\sum_{i=1}^{M}{\frac{v_{i}}{M}}$.
-In addition to the client’s values, we assume the privacy parameters ($\epsilon,\Delta$) and a shared random $seed$ are also provided.
-Clients Send Data
+These clients aim to collectively compute the element-wise sum of their private matrices: $\bm{v}=\sum_{i=1}^{M}{\frac{v_{i}}{M}}$.
+In addition to the client’s values, we assume the noise scale $\sigma$ and a random $seed$ (that all clients know) are also provided.
+Clients Send Data.
 The first step in the protocol is for each of the $M$ clients to generate a random mask set.
@@
 Because all the clients have access to a shared seed and $PRNG$ they can each compute the entire set locally.
-In Line 2, each client samples the entire set and sums it (in Line 2) to get the global mask matrix $\boldsymbol{r}$ (which will be used to decrypt later).
-In Line 2, each private value is converted to a fixed-point format, $\tilde{v_{i}}$, by scaling it up with a power-of-2 scale factor, $\mathbf{2^{\delta}}$.
-Formally, $\tilde{v_{i}}=\lceil{v_{i}\times\mathbf{2^{\delta}}}\rfloor$.
+In Line 1, each client samples the entire set and sums it (in Line 1) to get the global mask matrix $\bm{r}$ (which will be used to decrypt later).
+In Line 1, each private value is converted to a fixed-point format, $\tilde{v_{i}}$, by scaling it up with a power-of-2 scale factor, $2^{q}$.
+Formally, $\tilde{v_{i}}=\lceil{v_{i}\times 2^{q}}\rfloor$.
 All fractional values are rounded to the nearest representable value in this fixed-point representation.
-The choice of the scale factor, $\mathbf{2^{\delta}}$, determines the precision of the representation [^1] A larger $\mathbf{2^{\delta}}$ allows for greater precision but also reduces the number of bits available for the integer part of the number, which might cause overflows.
+The choice of the scale factor, $2^{q}$, determines the precision of the representation [^1] A larger $2^{q}$ allows for greater precision but also reduces the number of bits available for the integer part of the number, which might cause overflows.
 A workaround is to increase the bit-width of the operations, which increases the computational load and the communication cost..
-In practice, we empirically choose $\delta=16$.
-Then, in Line 2, each client encrypts (masks) their input $\tilde{v_{i}}$ by adding their share of the mask $r_{i}$.
+In practice, we empirically choose $q=16$.
+Then, in Line 1, each client encrypts (masks) their input $\tilde{v_{i}}$ by adding their share of the mask $r_{i}$.
 Finally, each client sends their masked values to the server.
@@
 […]
-Figure 2. Global Masked Secure Aggregation Protocol with $M$ Clients.
-Server Aggregates Data
-Upon receiving the masked matrices, the server first sums over each client’s contribution in Line 2.
-This yields the masked global sum $\boldsymbol{v}+\boldsymbol{r{}}$.
-The server then samples from a Laplace distribution to add noise proportional to $\Delta/\epsilon$ for each entry in the result matrix.
-This noise must also be converted to fixed-point by computing $\tilde{\gamma}=\lceil{\gamma\times\mathbf{2^{\delta}}}\rfloor$.
-We show why this preserves differential privacy in Appendix A.
-Finally, the server adds the noise $\tilde{\gamma}$ to the sum and broadcasts it back to the clients.
-Client Unmasks Data
-Upon receiving the result back from the server, each client must unmask the result.
-They do this by simply subtracting the mask $r{}$ that they computed in Line 2.
+Figure 1: Global Masked Secure Aggregation Protocol with $M$ Clients.
+Server Aggregates Data.
+Upon receiving the masked matrices, the server first sums over each client’s contribution in Line 1.
+This yields the masked global sum $\bm{v}+\bm{r{}}$.
+The server then samples from a zero mean Gaussian distribution with standard deviation equal to the supplied $\sigma$ for each entry in the result matrix.
+This noise must also be converted to fixed-point by computing $\tilde{\gamma}=\lceil{\gamma\times 2^{q}}\rfloor$.
+We show why this preserves differential privacy in Appendix A.3.
+Finally, the server adds the noise $\tilde{\gamma}$ to the sum and broadcasts it to the clients.
+Client Unmasks Data.
+Upon receiving the result from the server, each client must unmask the result.
+They do this by simply subtracting the mask $r{}$ that they computed in Line 1.
 After unmasking, each client must scale down the result to retrieve the correct answer.
-This is done by dividing the unmasked sum by the scale factor, $\mathbf{2^{\delta}}$, to reverse the initial scaling operation.
-5.2.5. Postprocessing
-In Algorithm 1, each client locally postprocesses the results (since adding noise to the centroids can result in values outside of the domain).
-A naive postprocessing strategy is to simply truncate out-of-bounds centroids to the boundary.
-However, in practice, we find that folding (Holohan et al., 2019) the value (reflecting it over the boundary) gives better utility.
-More formally, if $x<-B$, we adjust it to $x^{\prime}=-2B-x$, and if $x>B$, we adjust it to $x^{\prime}=2B-x$.
-This folding process effectively “folds” (mirrors) the value over the nearest boundary, preserving its distance from the boundary.
-We show in Appendix B that folding improves the utility of the algorithm, especially when the noise is large (small privacy budget).
-5.3. Proof of Privacy
+This is done by dividing the unmasked sum by the scale factor, $2^{q}$, to reverse the initial scaling operation.
+5.2.5 Post-processing
+In Algorithm 1, each client locally post-processes the results to obtain the centroids for the next iteration.
+First, each party divides the relative sum by the count and shifts the result by the previous centroid to get the new centroid (Line 19) following Eqn 10.
+Then, if the new centroid is more than $\eta$ away from the previous centroid, we truncate it to be $\eta$ away.
+Finally, we apply a post-processing step to the centroids to ensure they remain within the domain.
+A naive post-processing strategy is to simply truncate out-of-bounds centroids to the boundary.
+However, in practice, we find (in Appendix B) that folding [32] the value (reflecting it over the boundary) gives better utility.
+More formally, the operation folds a value $x$ into the range $[-B,B]$ through modular arithmetic.
+We first compute $(x+B)\bmod(2B)$, then if this result exceeds $B$, we reflect it about $B$ by subtracting it from $2B$.
+Finally, we return the value to the target range by subtracting $B$.
+This approach creates a periodic folding pattern that naturally reflects values across the boundaries while preserving distances from the nearest boundary point.
+We show in Appendix B that folding improves the algorithm’s utility.
+5.3 Privacy Analysis
 FastLloyd allows for a central party to add noise, retaining the utility of the central model of DP.
 However, since the server only interacts with masked values, and the clients can only observe a differentially private view of the final (noised) centroids, the protocol provides a level of privacy akin to that of the local model of DP.
-We prove the end-to-end privacy of our algorithm in the following theorem.
+We state the end-to-end security Theorem of our algorithm and defer the proof to Appendix A.2.
 Theorem 5.2.
-Algorithm 1 ensures $\epsilon(\lambda)$-IND-CDP-MPC in the presence of a semi-honest, polynomial time adversary who controls at most a single party. Where
-$\epsilon=Td\epsilon^{\mu}$
+Algorithm 1 ensures $(\epsilon(\lambda),\delta)$-IND-CDP-MPC in the presence of a semi-honest, polynomial time adversary who controls at most a single party.
+5.4 Error Analysis
+To analyze the error of our approach, we follow a similar approximate error analysis as Su et al.
+[64].
+The purpose of the analysis is primarily to choose the ratio of the privacy parameters and the number of iterations.
+Thus, the analysis makes a series of approximations.
+Following Su et al., we consider a single iteration and a single cluster for this analysis.
+To simplify the notation, we omit the cluster index $j$ and instead index the variables by the dimension $h$.
+We analyze the mean-squared error between the true centroid ($\mu$) and the differentially-private centroid ($\tilde{\mu}$) for one iteration across all dimensions.
+$MSE(\tilde{\mu}^{(t)})=\mathbb{E}\left[\sum\limits_{h=1}^{d}(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)})^{2}\right]$ (15)
+We first expand the following term using the definition of our DP mechanism from Section 5.1:
+$\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)}$ $=$ $\frac{(C^{(t)}+\gamma^{C})\mu_{h}^{(t)}}{C^{(t)}+\gamma^{C}}-\frac{S_{h}^{(t)}+\gamma^{R}_{h}+\gamma^{C}\tilde{\mu}_{h}^{(t-1)}}{C^{(t)}+\gamma^{C}}$
+$=$ $\frac{\gamma^{C}(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t-1)})-\gamma^{R}_{h}}{C^{(t)}+\gamma^{C}}.$
+Let $\nabla_{h}$ = $\mu_{j}^{(t)}-\tilde{\mu}_{j}^{(t-1)}$. Then, after squaring, we get:
+$(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)})^{2}=\frac{(\gamma^{C})^{2}(\nabla_{h})^{2}-2\gamma^{C}\gamma^{R}_{h}\nabla_{h}+(\gamma^{R}_{h})^{2}}{(C^{(t)}+\gamma^{C})^{2}}.$
+After taking the expectation over all clusters and assuming that $C+\gamma^{C}\approx N/k$, following Su et al. [64], we get:
+$\mathsf{MSE}\left(\tilde{\mu}_{h}^{(t)}\right)$$\approx$ $\frac{k^{3}}{N^{2}}\left(\mathbb{E}\left[(\gamma^{C})^{2}\right]\mathbb{E}\left[\nabla_{h}\right]^{2}+\mathbb{E}\left[(\gamma^{R}_{h})^{2}\right]\right)$
+$=$ $\frac{k^{3}}{N^{2}}\left(\mathsf{Var}\left(\gamma^{C}\right)\mathbb{E}\left[\nabla_{h}\right]^{2}+\mathsf{Var}\left(\gamma^{R}_{h}\right)\right)$
+where, in the first line, the middle term goes is zero as the noise terms are i.i.d. and zero mean. The second line holds because each $\gamma$ is an independent variable with zero mean, and so
+$\mathsf{E}[(\gamma)^{2}]=\mathsf{Var}\left(\gamma\right)-(\mathsf{E}[\gamma])^{2}=\mathsf{Var}\left(\gamma\right).$
+Finally, we approximate $\mathsf{E}[\sum_{h=1}^{d}\nabla_{h}]\approx\frac{\eta}{2}$.
+We argue that if the data were uniformly distributed in a hypersphere of radius $\eta$ around the centroid, the expected distance from the centroid would be $\eta/2$.
+This gives the final error term over all dimensions:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{4N^{2}}\left(\mathsf{Var}\left(\gamma^{C}\right)\eta^{2}+4d\mathsf{Var}\left(\gamma^{R}\right)\right)$ (16)
+Privacy Budget Splitting
+We now use the approximate analysis in Eqn 16 to determine the optimal privacy budget split between the relative sum and the count.
+We substitute the variance of the noise terms from Eqn 7 into Eqn 16:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}\eta^{2}T}{4N^{2}}\left((\sigma^{C})^{2}+4d(\sigma^{R})^{2}\right).$ (17)
+In our privacy analysis, we compute the noise multiplier $\sigma$ such that we achieve $\frac{1}{\sigma}$-GDP.
+Thus, we need to split the noise multiplier between the relative sum and the count following Theorem 2.2:
+$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (18)
+We minimize the per iteration error, subject to this constraint, using Lagrange multipliers, which gives:
+$\sigma^{C}=\sqrt[4]{4d}\sigma^{R}.$ (19)
+Simply scaling each sigma by this ratio gives Eqn 7. Substituting Eqn 7 back into the error analysis gives:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)=\frac{k^{3}\eta^{2}T\sigma^{2}(1+\sqrt{4d})^{2}}{4N^{2}}$ (20)
+Optimal Number of Iterations
+Using this analysis, we can determine a heuristic for the number of iterations.
+Following Su et al.
+[64], we assume that the error in each iteration is less than $0.004\times B$.
+Re-arranging for $T$ gives:
+$T<\frac{4N^{2}(0.004)}{k^{3}\eta^{2}\sigma^{2}(1+\sqrt{4d})^{2}}.$ (21)
+However, as Su et al.
+[64] explain, in practice, there is no need to go beyond seven iterations and at least two iterations are needed to gain useful results.
+Therefore, we truncate this analysis such that the number of iterations is in the range $[2,7]$.
+5.5 Setting the Maximum Radius Parameter
+We derive a dimensionality-aware radius parameter $\eta$ based on the geometric properties of the feature space.
+Consider a $d{}$-dimensional bounded feature space where each dimension is constrained to the interval $[-B{},B{}]$.
+The domain diagonal $\beta$, representing the maximum possible distance between any two points in this space, is given by:
+$\beta=\sqrt{d{}}\times(2B{})$ (22)
+When partitioning this space into $k{}$ clusters, each cluster occupies a fraction of the total volume.
+The total volume of the feature space is $(2B{})^{d{}}$.
+Assuming uniform partitioning, each cluster occupies a volume of $(2B{})^{d{}}/k{}$.
+Consequently, the effective length of each cluster along any dimension can be expressed as:
+$L_{c}=\frac{2B{}}{k{}^{1/d{}}}$ (23)
+Assuming hypercube-shaped clusters, the maximum distance from a centroid to any point within the cluster (the cluster radius) is half the cluster’s diagonal. This can be calculated as:
+$\text{Cluster Radius}=\frac{L_{c}}{2}\times\sqrt{d{}}=\frac{B{}}{k{}^{1/d{}}}\times\sqrt{d{}}=\frac{\beta}{2k{}^{1/d{}}}$ (24)
+Based on this analysis, we propose a heuristic $\eta_{d{}}$ to constrain the cluster radius in a $d{}$-dimensional space as:
+$\eta_{d{}}=\frac{\alpha\beta}{2k{}^{1/d{}}}$ (25)
+where $\alpha$ is a scaling factor we set to $0.8$ based on experiments using synthetic data in Appendix B.
+In scenarios with low $d$ and high $k$, $\eta$ becomes increasingly restrictive as $k{}^{1/d{}}$ grows larger, providing tighter bounds on cluster radii.
+However, as dimensions increase, the curse of dimensionality necessitates larger cluster radii to accommodate the exponential growth of volume in the space, $(2B{})^{d{}}$.
+Our approach takes this into account through $k{}^{1/d{}}$ approaching unity in high dimensions.
+This dimensional scaling aligns with the intuition that radius constraints are most meaningful in lower-dimensional spaces, where cluster boundaries remain well-defined, whereas, in high-dimensional spaces, the curse of dimensionality renders such constraints increasingly less effective as distance metrics lose their discriminative power.
+We consider two distinct approaches to implement the radius constraint.
+In the “Constant” approach, the radius constraint is consistently enforced throughout the clustering process.
+Alternatively, the “Step” approach initializes with a broader constraint of $\beta/2$ at iteration zero and then transitions to the computed radius $\eta$ for the remaining iterations.
+This two-phase strategy allows for initial flexibility in centroid placement while gradually imposing stricter constraints for fine-tuning the centroids.
+We show in Appendix B that the “Step” approach is superior.
+6 Evaluation
+In this section, we provide an extensive evaluation of our work in terms of the following questions:
+  1. Q1
+    How does the utility of FastLloyd  compare with state-of-the-art in central model [64] for interactive DP $k$-means?
+  2. Q2
+    How does the utility of FastLloyd scale with varying the number of dimensions and number of clusters?
+  3. Q3
+    How does the runtime, communication, and number of rounds for FastLloyd compare with the state-of-the-art in federated $k$-means using secure computation [50]?
+  4. Q4
+    How does the runtime and communication of FastLloyd  scale with varying the number of dimensions, clusters, and data points?
+(a) Adult
+(b) Birch2
+(c) Iris
+(d) Breast
+(e) Yeast
+(f) House
+(g) LSun
+(h) S1
+(i) Wine
+(j) MNIST
+Figure 2: Normalized Intra-cluster Variance (NICV) vs $\epsilon$ for real datasets.
+6.1 Experimental Setup
+6.1.1 Implementation
+Our experimental evaluation was conducted on a Macbook Pro M2 Max (30-core CPU, 38-core GPU, 64GB RAM) using Open MPI [26] for multiparty communication.
+Following Mohassel et al.
+[50], we evaluate in the LAN setting with simulated network latency (0.25ms per send operation), noting that WAN runtimes can be derived using a linear cost model.
+Our experimental framework utilizes a two-client setup with balanced dataset partitioning.
+The MSA protocol (Figure 1) operates over the ring $\mathbf{R}=\mathcal{Z}_{2^{32}}$, employing fixed-point representation with precision factor $q=16$.
+All experiments use randomly partitioned datasets, with results averaged over 100 runs and reported with 95% confidence intervals where applicable.
+For FastLloyd, we set the radius constraint $\alpha=0.8$ using the “Step” strategy (see Appendix B for detailed analysis and justification of these parameter choices).
+We set $\delta=\frac{1}{N\log{N}}$ for all experiments.
+Our implementation is publicly available at https://github.com/D-Diaa/FastLloyd.
+6.1.2 Baselines
+We evaluate FastLloyd against several baselines.
+For evaluating utility, we implement three protocols.
+First, Lloyd [44], a non-private Lloyd’s algorithm adapted to our federated framework with sphere packing initialization.
+Second, SuLloyd [64], which adapts Su et al.’s centralized differentially private $k$-means clustering algorithm to our federated setting using our masked secure aggregation (MSA).
+Third, GLloyd, a modification of SuLloyd replacing Laplace noise with Gaussian noise and using composition and privacy budget analysis similar to our work.
+In Appendix B.1, we provide the details of how the analysis in Section 5.4 differs for GLloyd.
+For computational and communication efficiency benchmarking, we compare against MohLloyd [50], established by Hedge et al.
+[31] ass the most efficient secure $k$-means protocol.
+6.1.3 Datasets
+Our evaluation employs both real and synthetic datasets.
+We use established datasets from the clustering datasets repository [25], following Su et al.
+[64] and Mohassel et al.
+[50] for comparability.
+For Birch2 [72], we take $25,000$ random samples from the $100,000$ sample dataset.
+For evaluating scalability, following [64], we generate synthetic datasets (Synth) using the clusterGeneration R package [58], which enables control over inter-cluster separation (in $[-1,1]$).
+The synthetic datasets contain $N=10,000$ samples across $k$ clusters, with cluster sizes following a $1:2:...:k$ ratio.
+We incorporate a random number (in $[0,100]$) of randomly sampled outliers and set the cluster separation degrees in $[0.16,0.26]$, spanning partially overlapping to separated clusters.
+While Su et al.
+[64] evaluate configurations up to $k=10$ and $d=10$, we extend the evaluation to $k=32$ and $d=512$ in powers of two, creating $45$ parameter combinations.
+Each combination generates three datasets with different random seeds.
+To assess scalability at higher cluster counts, we further extend our evaluation with Synth-K, incorporating configurations up to $k=128$ for $d=2$.
+For benchmarking performance, we create synthetic datasets (TimeSynth) with balanced cluster sizes ($C_{avg}=\frac{N}{k}$), varying $N$ (10K-100K), $d$ (2-5), and $k$ (2-5), following Mohassel et al.
+[50].
+Table 2 summarizes these datasets.
+Following standard practice in DP literature [9, 47, 16, 64], all datasets are normalized to $[-1,1]$.
+Table 2: Summary of Datasets Used in Evaluation
+Dataset $N$ $d$ $k$
+Iris [23] 150 4 3
+LSun [65] 400 2 3
+S1 [24] 5000 2 15
+House [25] 1837 3 3
+Adult [7] 48842 6 3
+Wine [1] 178 13 3
+Breast [67] 699 9 2
+Yeast [51] 1484 8 10
+MNIST [41] 60000 784 10
+Birch2 [72] 25000 2 100
+G2 [46] 2048 2-1024 2
+Synth 10K 2-512 2-32
+Synth-K 10K 2 2-128
+TimeSynth 10K, 100K 2,5 2,5
+(a) Varying the number of dimensions using G2 ($k=2$, log scale).
+(b) Varying the number of clusters using Synth-K ($d=2$).
+Figure 3: Scaling of AUC over number of dimensions and clusters.
+(a) SuLloyd
+(b) FastLloyd
+(c) Improvement
+Figure 4: AUC comparison of FastLloyd vs. SuLloyd on the Synth datasets.
+6.2 Utility Evaluation
+6.2.1 Metrics
+We use the Normalized Intra-cluster Variance (NICV) from prior work [64] as our primary utility metric. NICV normalizes the $k$-means objective function (Eqn 1) by dividing it by the dataset size:
+$NICV(O)=\frac{1}{N}\cdot\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
+To facilitate systematic comparison across methods for the synthetic datasets, we summarize each method’s performance over different privacy budgets using the area under the curve (AUC) of NICV values against $\epsilon\in[0.1,0.25,0.5,0.75,1.0]$, computed via the trapezoidal rule:
+$AUC=\sum_{i=1}^{n-1}\frac{NICV_{i}+NICV_{i+1}}{2}\cdot(\epsilon_{i+1}-\epsilon_{i})$ (26)
+6.2.2 Real World Datasets
+Figure 2 shows the NICV against various privacy budgets on ten real-world datasets to answer Q1.
+The shading shows the 95% confidence interval of the mean over the 100 experiments.
+We observe that FastLloyd and GLloyd consistently outperform SuLloyd across all high-dimensional datasets (e.g., Wine, Breast, Yeast, MNIST).
+This aligns with our theoretical expectations as the Gaussian mechanism composes much more favourably in high-dimensional spaces than the Laplace mechanism.
+Additionally, FastLloyd outperforms GLloyd in all cases, especially in datasets with low dimensions and a high number of clusters (e.g., Birch2, S1), where GLloyd does not outperform SuLloyd.
+This highlights the effect of imposing a radius constraint on the protocol, which, as discussed in Section 5.1, is most restrictive in low-dimensional, high-cluster settings, further reducing the noise added to the centroids.
+While all differentially private methods eventually approach the non-private baseline as privacy budget increases, FastLloyd achieves this convergence at substantially lower privacy budgets, establishing it as the superior approach for DP $k$-means clustering across all datasets.
+6.2.3 Synthetic Datasets
+To answer Q2, we follow an approach similar to Su et al.
+[64] and evaluate FastLloyd and SuLloyd on our synthetic datasets varying both $k$ and $d$.
+Figure 4 shows heatmaps of the AUC values for SuLloyd and FastLloyd on the synthetic datasets.
+To clarify the difference between the protocols, Figure 4(c) shows the percentage of reduction in AUC of FastLloyd over that of SuLloyd.
+From the heatmaps, we observe that FastLloyd outperforms SuLloyd across all datasets, with the improvement being more pronounced in higher dimensions and in higher $k$.
+To investigate higher limits, we further extend the evaluation to the G2 [46] datasets with $d$ up to 1024, and Synth-K datasets with $k$ up to 128.
+Figures 3(a) and 3(b) present a comparative analysis of AUC performance across Lloyd, FastLloyd, GLloyd, and SuLloyd protocols, evaluated on G2 and Synth-K datasets.
+The AUC values are normalized relative to SuLloyd as the baseline.
+Figure 3(a) examines dimensionality scaling by varying $d$ up to 1024 while maintaining $k=2$, whereas Figure 3(b) shows cluster scaling by varying $k$ up to 128 while fixing $d=2$.
+In Figure 3(a), we observe that for low dimensions, all protocols perform very closely to the baseline (Lloyd), but as the dimension increases, the task becomes much more challenging for the private protocols, with FastLloyd outperforming all others.
+In Figure 3(b), we observe that FastLloyd outperforms all other protocols across all cluster counts, with the improvement being more pronounced at higher cluster counts.
+This is to be expected as the radius constraint becomes more restrictive with higher cluster counts, leading to a more significant reduction in noise added to the centroids.
+Since Su et al.
+found SuLloyd outperformed all other approaches for $d{}>3$ [64], and we outperform SuLloyd, we claim that FastLloyd is state-of-the-art in DP $k$-means for $d{}>3$.
+FastLloyd also fixes the scalability issue that Su et al.’s work [64] observed in high $k$ and enables us to scale to much larger dimensions.
+Dataset Parameters Runtime (ms) Comm (bytes)
+$N$ $k$ $d$ FastLloyd MohLloyd [50] Speedup FastLloyd MohLloyd [50] Reduction
+TimeSynth 10K 2 2 $3.03\pm 0.02$ 10500 3465×3465$\times$3465 × 192 2.557e8 $1.33e6\times$
+2 5 $3.24\pm 0.01$ 384 - -
+5 2 $4.32\pm 0.02$ 34050 7882×7882$\times$7882 × 480 9.746e8 $2.03e6\times$
+5 5 $4.95\pm 0.04$ 960 - -
+TimeSynth 100K 2 2 $12.57\pm 0.02$ 105120 8363×8363$\times$8363 × 192 2.467e9 $1.28e7\times$
+2 5 $13.17\pm 0.03$ 384 - -
+5 2 $22.05\pm 0.03$ 347250 15748×15748$\times$15748 × 480 9.535e9 $1.99e7\times$
+5 5 $22.56\pm 0.02$ 960 - -
+LSun 400 3 2 $2.6\pm 0.04$ 1481 570×570$\times$570 × 288 - -
+S1 5K 15 2 $5.81\pm 0.08$ 49087 8449×8449$\times$8449 × 1440 - -
+Table 3: Overhead comparison per iteration against MohLloyd [50] for two clients ($100$ runs with mean and $95\%$ confidence reported for FastLloyd)
+6.3 Runtime Evaluation
+To answer questions Q3 and Q4, we compare FastLloyd with the MohLloyd protocol of Mohassel et al.
+[50] on a variety of common datasets in terms of runtime and communication size (per iteration).
+Table 3 shows the summary of the evaluation with two clients (the setting considered in Mohassel et al.’s work [50]).
+Values for MohLloyd are taken as reported in the paper by Mohassel et al.
+[50], noting that the setup is similar to the one used in our evaluation, and the gap in performance is more than what could be accounted for by different setups.
+In terms of runtime, MohLloyd executes in the order of minutes, while FastLloyd executes in milliseconds, offering five orders of magnitude speedup.
+In terms of communication size, MohLloyd requires communicating gigabytes of data per iteration, while FastLloyd needs a fraction of a kilobyte, offering up to seven orders of magnitude reduction in size.
+In terms of communication rounds, MohLloyd requires $\Theta(\lceil\log{k{}}\rceil)$ communication rounds per iteration [31], while FastLloyd only requires one.
+Since MohLloyd was found to be the state-of-the-art in secure federated $k$-means [31], we conclude that we advance the state-of-the-art while offering output privacy (which would only further slow down Mohassel et al.
+[50]).
+7 Conclusion
+In this work, we designed FastLloyd, a new private protocol for federated $k$-means.
+We have shown that FastLloyd is secure in the computational model of DP and analyzed its utility.
+Compared to state-of-the-art solutions in the central model of DP, FastLloyd results in higher utility across a wide range of real datasets and scales effectively to larger dimensions and number of clusters.
+FastLloyd also achieves five orders of magnitude faster runtime than the state-of-the-art in secure federated $k$-means across a variety of problem sizes.
+In summary, we provide an efficient, private, and accurate solution to the horizontally federated $k$-means problem.
+Acknowledgments
+We gratefully acknowledge the support of NSERC for grants RGPIN-2023-03244, IRC-537591, the Government of Ontario and the Royal Bank of Canada for funding this research.
+Appendix A Proofs
+A.1 Proof of Theorem 5.1
+See 5.1
 Proof.
-In order to prove the algorithm satisfies Definition 3.1, we need to consider the view from each party.
+w.l.o.g assume that the datasets differ by a single point $x^{\prime}$ so that $D^{\prime}=D\cup\{x^{\prime}\}$.
+If $x^{\prime}$ is not within $\eta$ of $\tilde{\mu}_{j}^{(t-1)}$, it will not be assigned to $O_{j}$ and thus $R_{j}(D)=R_{j}(D^{\prime})$.
+Therefore, we only need to consider the case where $x^{\prime}$ is at most $\eta$ from $\tilde{\mu}_{j}^{(t-1)}$.
+By definition, we have:
+$\Delta^{R}$ $=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}-(S_{j}^{\prime(t)}-C_{{}^{\prime}}{j}^{(t)}\tilde{\mu}_{j}^{(t-1)})||_{2}$
+$=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}-S_{j}^{(t)}-x^{\prime}+(C_{j}^{(t)}+1)\tilde{\mu}_{j}^{(t-1)}||_{2}$
+$=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||\tilde{\mu}_{j}^{(t-1)}-x^{\prime}||_{2}$
+$\leq$ $\eta$
+where the last line follows by (Eqn 4). ∎
+A.2 Proof of Theorem 5.2
+We first prove the following helpful lemma.
+Lemma A.1.
+Our noise mechanism (defined in Eqn 8) of $f(D)+\gamma$ where $\gamma\sim\mathcal{N}(0,\sigma^{2}T(\Delta^{(f)})^{2})$ applied over $T$ adaptive iterations of $f$ is $(\frac{1}{\sigma})$-GDP, where $\Delta^{(f)}$ is the sensitivity of $f$.
+Proof.
+After choosing $\sigma$, by Theorem 2.1 from Dong et al.
+[15, Theorem 1], we have that each application of a Gaussian mechanism with noise multiplier $\sigma$ is $\frac{\Delta^{(f)}}{\sigma}-GDP$.
+We apply the Gaussian mechanism adaptively over $T$ iterations.
+Which by Theorem 2.2 from Dong et al.
+[15, Corollary 2], gives us $\frac{\Delta^{(f)}\sqrt{T}}{\sigma}$-GDP.
+Thus, by multiplying $\sigma$ by $\sqrt{T}\Delta^{(f)}$, we get $\frac{1}{\sigma}$-GDP.
+∎
+We now restate and prove Theorem 5.2. See 5.2
+Proof.
+To prove the algorithm satisfies Definition 3.1, we need to consider the view from each party.
 We begin with the view of the server.
@@
 Therefore, the view of the server satisfies $\epsilon(\lambda)$-IND-CDP-MPC with $\epsilon=0$.
-In regards to the clients, we assume no client colludes with any other client.
-Then, a client’s view consists of their own dataset $D_{A}$, and the output from each iteration of the protocol.
+Regarding the clients, a client’s view consists of their own dataset $D_{A}$, and the output from each iteration of the protocol.
 We note that the initialization is data-independent and thus is indistinguishable between neighbouring datasets.
 Each iteration consists of the assignment and updating of the cluster centroids.
-The assignment step is simply postprocessing on the previous iteration (or random initialization) and thus incurs no additional privacy cost.
-The updating of the cluster centroids applies the Laplace mechanism to each dimension separately, with the sensitivity parameter set following Eqn 6 and $\epsilon^{\mu}$ as the privacy budget.
-Using naive, sequential composition over the dimensions and the iterations, the client’s view is $Td\epsilon^{\mu}$-IND-CDP-MPC, with no computational assumption (as we use the information-theoretic DP properties of the Laplace mechanism and composition (Dwork and Roth, 2014)).
+The assignment step uses the published centroids from the previous iteration (post-processing) to divide the dataset into clusters.
+We can then apply parallel composition over each of the clusters.
+Thus, we can focus on the privacy cost of a single cluster for the remainder of the proof.
+First, we set $\sigma$ using Theorem 2.3 from Dong et al.
+[15, Corollary 1].
+We choose the minimum $\sigma$ such that $(\epsilon,\delta)$-DP iff $\frac{1}{\sigma}$-GDP by Theorem 2.3.
+This is equivalent to finding the noise multiplier for a sensitivity one, single application, of the Gaussian mechanism with noise multiplier $\sigma$.
+To solve this minimization, we use Algorithm 1 from Balle and Wang [6].
+The updating of the cluster centroid applies the Gaussian mechanism twice, once to the relative sums and once to the counts.
+We split the overall $\sigma$ into $\sigma^{R}$ and $\sigma^{C}$ following Eqn 7.
+We begin with the analysis of the relative sum.
+In Theorem 5.1, we show that the sensitivity of the relative sum is $\eta$.
+Applying Lemma A.1 we get that the relative sum over $T$ iterations is $\frac{1}{\sigma^{R}}$-GDP.
+The sensitivity of the count is $1$, and thus the count is $\frac{1}{\sigma^{C}}$-GDP by similar analysis.
+Finally, applying Theorem 2.2.
+We get that
+$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (27)
+and thus the we get $\frac{1}{\sigma}$-GDP over the entire protocol which implies $(\epsilon,\delta)$-DP (because of how we chose $\sigma$).
+Applying parallel composition over all clusters, the client’s view is $(\epsilon,\delta)$-IND-CDP-MPC, with no computational assumption (as we use the information-theoretic DP properties of differential privacy).
 Taking the worst case over the clients and the server, the result follows.
 ∎
-5.4. Error Analysis
-To analyze the error of our approach, we follow a similar analysis to Su et al.
-(Su et al., 2016).
-We analyze the mean-squared error between the true centroid ($\mu_{j}$) and the differentially-private centroid ($\bar{\mu_{j}}$) for one iteration across all dimensions.
-Substituting the definition of the centroid method (Eqn 7), we get
-$\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=\mathsf{E}\left[\sum_{h=1}^{d}((\mu_{j}[h]+\gamma^{\mu})-\mu_{j}[h])^{2}\right]=$
-$=\sum_{h=1}^{d}\mathsf{E}\left[(\gamma^{\mu})^{2}\right]=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right).$
-Where the last line holds because $\gamma^{\mu}$ is an independent Laplacian variable with zero mean and so
-$\mathsf{E}[(\gamma^{\mu})^{2}]=\mathsf{Var}\left(\gamma^{\mu}\right)-(\mathsf{E}[\gamma^{\mu}])^{2}=\mathsf{Var}\left(\gamma^{\mu}\right).$
-Plugging in the sensitivity of the mean (Eqn 6) and the privacy budget gives
-(9) $\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right)=\frac{2dB^{2}}{(\epsilon^{\mu}\eta_{min})^{2}}$
-(10) $=\frac{2d^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}=\mathcal{O}\left(\frac{d^{3}T^{2}}{\epsilon^{2}\eta_{min}^{2}}\right).$
-This shows that the distortion in the centroid is proportional to $d^{3}T^{2}$ and inversely proportional to $\epsilon^{2}\eta_{min}^{2}$.
-Remark.
-This error is similar to what Su et al.
-(Su et al., 2016) observed for the sum-count method with the following differences: (1) $\eta_{min}$replaces the assumed expected cluster size of $\frac{N}{k}$.
-(2) The error is independent of the data or the actual clustering result; it depends completely on the parameters set before running the algorithm.
-(3) This is a worst-case analysis, and the actual error might be lower.
-Adding across all clusters gives
-(11) $\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}.$
-Using this analysis, we can determine the optimal number of iterations, following a similar approach as Su et al.
-(Su et al., 2016).
-First, we calculate a minimum privacy budget per iteration ($\epsilon^{(m)}$) so that the mean-squared error is within a certain bound.
-Secondly, we divide the available budget ($\epsilon$) by $\epsilon^{(m)}$ in order to determine the number of iterations $T$.
-Thus, setting $T=1$ and $\epsilon=\epsilon^{(m)}$ gives us
-$\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}B^{2}}{(\epsilon^{(m)}\eta_{min})^{2}}\leq 0.004$
-which implies
-$\sqrt{\frac{500kd^{3}B^{2}}{\eta_{min}^{2}}}\leq\epsilon^{(m)}.$
-However, Su et al.
-note that, in practice, there is no need to go beyond seven iterations and that at least two iterations are needed to gain useful results.
-Therefore, we truncate the number of iterations to be in the range $[2,7]$.
-6. Evaluation
-In this section, we provide an extensive evaluation of our work in terms of the following questions:
-  1. Q1
-    How does the utility of FastLloyd  compare with state-of-the-art in central model (Su et al., 2016) for DP $k$-means?
-  2. Q2
-    How does constrained clustering affect the utility of the sum-count method?
-  3. Q3
-    How does the utility of FastLloyd scale with varying the number of dimensions and number of clusters?
-  4. Q4
-    How does the runtime, communication, and number of rounds for FastLloyd compare with the state-of-the-art in federated $k$-means using secure computation (Mohassel et al., 2020)?
-  5. Q5
-    How does the runtime and communication of FastLloyd  scale with varying the number of dimensions, clusters, and data points?
-6.1. Experimental Setup
-(a) S1
-(b) Lsun
-(c) Birch2
-(d) House
-(e) Iris
-(f) Adult
-Figure 3. Utility evaluation of FastLloyd vs. SuLloyd on real datasets (20 runs).
-Implementaion
-The experiments were carried out on a Macbook Pro M2 Max, with a 30-core CPU, a 38-core GPU, and 64 GB of RAM.
-We use the Open MPI (Gabriel et al., 2004) (Open Source High-Performance Computing) library to implement multiparty communication through message-passing.
-All experiments were conducted in the multiparty framework, with datasets being randomly split between parties.
-Following Mohassel et al.
-(Mohassel et al., 2020), we only evaluate in the LAN setting, and observe that WAN runtimes can be computed with the linear cost model (adding a fixed latency for each round).
-We simulate the network latency by introducing a custom (sleep) delay of 0.5ms for every “send” operation.
-We make our code publicly available to reproduce the results222 https://github.com/D-Diaa/privateclustering.git.
-All results were averaged over 20 runs and the 95% mean confidence interval is supplied where appropriate.
-Baselines
-To evaluate the clustering quality of FastLloyd, we compare it against the state-of-the-art in central DP for $k$-means clustering by Su et al.
-(Su et al., 2016), denoted as SuLloyd.
-As established in Section 4, SuLloyd is the most accurate DP $k$-means clustering algorithm to date.
-We implement a federated version of SuLloyd in our framework for fair evaluation.
-To study the effect of constrained clustering on the sum-count method used in SuLloyd, we also extend the federated implementation of SuLloyd to use the same constraints as FastLloyd, denoted as CSuLloyd.
-Finally, we give a non-private baseline of the original Lloyd’s algorithm (Lloyd, 1982), implemented into our federated framework (with sphere packing initialization) and denoted as Lloyd.
-To evaluate the computational and communicational costs in the secure federated setting, we compare against the corresponding state-of-the-art by Mohassel et.
-al (Mohassel et al., 2020), denoted as MohLloyd.
-According to the SoK of Hedge et al.
-(Hegde et al., [n.
-d.]), MohLloyd is the most efficient secure $k$-means clustering algorithm to date.
-For MohLloyd, we evaluate against the reported values in the original paper (Mohassel et al., 2020).
-Parameters
-In our experiments, we use the following default parameters, unless otherwise specified.
-The number of clients is set to two, and the privacy budget $\epsilon=1$.
-For the MSA protocol (Figure 2), the underlying ring $\mathbf{R}$ is $\mathcal{Z}_{2^{32}}$ and the values are all transformed to fixed-point under precision factor $\delta=16$.
-For any dataset of size $N{}$, to be clustered into $k{}$ clusters, we fix $\eta_{min}=\frac{N{}}{1.25k{}}$ and $\eta_{max}=\frac{1.25N{}}{k{}}$ (empirically chosen as per Appendix B).
-Datasets
-We selected a wide range of commonly used datasets to evaluate our approach, we follow Su et al.
-(Su et al., 2016) and Mohassel et al.
-(Mohassel et al., 2020) to maximize comparability.
-An exception is the Birch2 (Zhang et al., 1997) dataset, which is a synthetic dataset comprised of 100 Gaussian clusters forming a sine wave.
-This dataset initially contains 100K data points, however, a random sample of 25K points is used.
-To the best of our knowledge, we are the first federated privacy-preserving work to evaluate using a dataset of this size (in terms of $N$ and $k$).
-Additionally, we created synthetic datasets to evaluate the scalability of our approach.
-Similar to Su et al.
-(Su et al., 2016), we use the R package (Qiu and Joe, 2023), clusterGeneration.
-The package is designed for generating cluster datasets with a given degree of separation ($\in[-1,1]$) between neighbouring clusters.
-For the clustering quality evaluation, we generate two sets of datasets: one with overlapping cluster structure (SynthOV, separation = 0.21) and one with well-separated cluster structure (SynthSEP, separation = 0.70).
-Each set varies $k$ and $d$ from $2$ to $20$ in steps of $2$, resulting in 100 variations.
-For every variation, we generate three different datasets with different seeds.
-For all synthetic datasets, we fix an average cluster size of $C_{avg}=\frac{10K}{k}$ and randomly sample cluster sizes to be in the range $[0.75\cdot C_{avg},1.25\cdot C_{avg}]$, resulting in a dataset that has a total size in the range $[7.5K,12.5K]$.
-For our timing evaluation, we generate similar synthetic datasets, TimeSynth, similar to those in Mohassel et al.’s work (Mohassel et al., 2020), enforcing an exact cluster size of $C_{avg}=\frac{N}{k}$.
-We vary $N$ between 10K and 100K, $d$ between two and five and $k$ between two and five.
-A summary of these datasets is presented in Table 2.
-As is common in the DP literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), we assume all datasets are normalized in the range $[-1,1]$ as pre-processing.
-Table 2. Summary of Datasets Used in Evaluation
-Dataset $N$ $d$ $k$
-Iris (Fisher, 1988) 150 4 3
-LSun (Ultsch, 2005) 400 2 3
-S1 (Fränti and Virmajoki, 2006) 5000 2 15
-House (Fränti and Sieranoja, 2018) 1837 3 3
-Adult (Becker and Kohavi, 1996) 48842 6 3
-Birch2 (Zhang et al., 1997) 25000 2 100
-SynthOV 7.5K ~ 12.5K 2-20 2-20
-SynthSEP 7.5K ~ 12.5K 2-20 2-20
-TimeSynth 10K, 100K 2,5 2,5
-6.2. Utility Evaluation
-In this section, we focus on evaluating the utility of FastLloyd compared to the state-of-the-art work of Su et al. (Su et al., 2016). We begin by describing the metrics used for this evaluation.
-Metrics
-Following similar literature (Su et al., 2016), we use the Normalized Intra-cluster Variance (NICV) metric to evaluate clustering quality.
-In short, NICV is the $k$-means objective function (1) divided by the number of points in the dataset.
-$NICV(O)=\frac{1}{N}\cdot\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
-We also introduce another metric, Empty Clusters, that measures the ratio of empty clusters to $k$.
-The objective of $k$-means clustering is to produce exactly $k$ clusters such that the WCSS (NICV) is minimized.
-However, existing metrics/objectives do not capture the requirement of having exactly $k$ clusters.
-This becomes problematic in differentially private settings, where adding noise can isolate some centroids from the data completely, resulting in a suboptimal solution that uses less than $k$ clusters.
-We introduce this metric to precisely quantify this effect as follows:
-Empty $\texttt{Empty Clusters}(O)=\frac{\sum_{j=1}^{k}{I(|O_{j}|=0)}}{k}$
-6.2.1. Real World Datasets
-To answer Q1 and Q2, we evaluate FastLloyd, CSuLloyd, and SuLloyd on a variety of datasets.
-Figure 3 shows the NICV against various privacy budgets ($\epsilon$) from 0.01 to 2.0 on six datasets.
-The shading shows the 95% confidence interval over the 20 experiments.
-First, we consider Q1.
-We observe that FastLloyd outperforms SuLloyd on all datasets, especially for lower values of epsilon.
-On a larger dataset (birch2), Figure 3(c) shows a statistically significant improvement (the confidence intervals do not overlap).
-FastLloyd even outperforms the non-private baseline of Lloyd at $\epsilon=2$.
-For higher dimensional datasets, a substantial performance gap is evident; Figure 3(f) shows that even in high-privacy settings (with $\epsilon=0.01$), FastLloyd significantly outperforms the non-private baseline of Lloyd.
-To answer Q2, we observe that for all datasets CSuLloyd  outperforms SuLloyd.
-However, in all cases, FastLloyd outperforms SuLloyd (except in Figure 3(f) , where we tie).
-To evaluate the Empty Clusters metric, we use the birch2 dataset with a relatively high cluster count of 100.
-Figure 4 shows that even the non-private Lloyd loses around $15\%$ of clusters, with SuLloyd reaching up to more than $40\%$ in high-privacy settings (with $\epsilon=0.01$).
-We observe that FastLloyd eliminates this problem completely for $\epsilon>0.75$ and significantly diminishes it for smaller privacy budgets.
-Figure 4. Birch2 Empty Clusters
-(a) SuLloyd
-(b) FastLloyd
-(c) Improvement Percentage
-Figure 5. NICV Utility comparison over SynthOV dataset. Averaged over 60 runs (20 on each of three random datasets).
-(a) SuLloyd
-(b) FastLloyd
-(c) Improvement Percentage
-Figure 6. NICV Utility comparison over SynthSEP dataset. Averaged over 60 runs (20 on each of three random datasets).
-6.2.2. Synthetic Datasets
-To answer Q3, we follow an approach similar to Su et al.
-(Su et al., 2016) and evaluate FastLloyd and SuLloyd on SynthSEP and SynthOV, varying both $k$ and $d$.
-Figures 5 and 6 show heatmaps of NICV values for both approaches when $\epsilon=1$.
-To clarify the difference between the protocols, Figures 5(c) and 6(c) show the percentage of reduction in NICV of FastLloyd over that of SuLloyd.
-For the overlapping dataset (SynthOV), FastLloyd  outperforms SuLloyd in all cases, with the improvement being more significant in higher cluster counts and low dimensions, reducing NICV by more than 40%.
-The gap in performance becomes more pronounced in the separated dataset (SynthSEP), with FastLloyd  significantly outperforming SuLloyd in all cases, reducing the NICV error by up to 89% in some cases, and by more than 78% even in the highest number of clusters and dimensions.
-These observations highlight the better scalability and utility of FastLloyd over SuLloyd in high-dimensional and high-cluster-count scenarios.
-Since SuLloyd was found to outperform all other approaches for $d{}>3$ (Su et al., 2016), we claim that FastLloyd is state-of-the-art in DP $k$-means for $d{}>3$.
-FastLloyd also fixes the scalability issue that Su et al.’s work (Su et al., 2016) observed in high $k$.
-6.3. Runtime Evaluation
-Dataset Parameters Runtime (ms) Comm (bytes)
-$N$ $k$ $d$ FastLloyd MohLloyd (Mohassel et al., 2020) Speedup FastLloyd MohLloyd (Mohassel et al., 2020) Reduction
-TimeSynth 10K 2 2 $3.23\pm 0.23$ 10500 3250×3250$\times$3250 × 128 2.557e8 1,998,000×1,998,000$\times$1 , 998 , 000 ×
-2 5 $8.61\pm 1.81$ 320 - -
-5 2 $18.02\pm 3.59$ 34050 1890×1890$\times$1890 × 320 9.746e8 3,045,000×3,045,000$\times$3 , 045 , 000 ×
-5 5 $19.83\pm 2.01$ 800 - -
-TimeSynth 100K 2 2 $85.57\pm 2.42$ 105120 1228×1228$\times$1228 × 128 2.467e9 19,277,000×19,277,000$\times$19 , 277 , 000 ×
-2 5 $313.47\pm 6.79$ 320 - -
-5 2 $423.1\pm 6.88$ 347250 820×820$\times$820 × 320 9.535e9 29,797,000×29,797,000$\times$29 , 797 , 000 ×
-5 5 $412.89\pm 5.84$ 800 - -
-LSun 400 3 2 $1.81\pm 0.1$ 1481 1156×1156$\times$1156 × 192 - -
-S1 5K 15 2 $24.76\pm 0.3$ 49087 1982×1982$\times$1982 × 960 - -
-Table 3. Overhead comparison per iteration against MohLloyd (Mohassel et al., 2020) for two clients ($20$ runs with mean and $95\%$ confidence reported for FastLloyd)
-To answer questions Q4 and Q5, we compare FastLloyd with the MohLloyd protocol of Mohassel et al.
-(Mohassel et al., 2020) on a variety of common datasets in terms of runtime and communication size (per iteration).
-Table 3 shows the summary of the evaluation with two clients (the setting considered in Mohassel et al.’s work (Mohassel et al., 2020)).
-Values for MohLloyd are taken as reported in the paper by Mohassel et al.
-(Mohassel et al., 2020), noting that the setup is similar to the one used in our evaluation, and the gap in performance is more than what could be accounted for by different setups.
-In terms of runtime, MohLloyd executes in the order of minutes, while FastLloyd executes in a fraction of a second, offering four orders of magnitude speedup.
-In terms of communication size, MohLloyd requires communicating gigabytes of data per iteration, while FastLloyd needs a fraction of a kilobyte, offering up to seven orders of magnitude reduction in size.
-In terms of communication rounds, MohLloyd requires $\Theta(\lceil\log{k{}}\rceil)$ communication rounds per iteration (Hegde et al., [n.
-d.]), while FastLloyd only requires one.
-Since MohLloyd was found to be the state-of-the-art in secure federated $k$-means (Hegde et al., [n.
-d.]), we conclude that we advance the state-of-the-art while offering output privacy (which would only further slow down Mohassel et al.
-(Mohassel et al., 2020)).
-7. Extensions
-7.1. Distance-Constrained $k$-means
-(a) S1 NICV
-(b) Birch2 NICV
-(c) Birch2 Empty Clusters
-Figure 7. The effect of the distance constraint (dashed lines) on the clustering quality of FastLloyd and CSuLloyd
-[…]
-Figure 8. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$, $\boldsymbol{\eta}^{max}$ and $\eta^{dist}$.
-We introduce a new extension of the constrained clustering problem that further tunes and improves the utility of DP clustering.
-This new problem includes a constraint on the maximum radius ($\eta_{d}$) of every cluster, such that the objective function is now subject to the constraints:
-(12) $\eta_{min}\leq|O_{j}|\leq\eta_{max}\quad\text{and}\quad||x_{l}-\mu_{j}||\leq\eta_{d}\quad\forall x_{l}\in O_{j},\forall j$
-In lieu of previous work, we extend the network flow formulation to solve our new constrained clustering problem; an example is shown in Figure 8.
-In short, we modify the capacity of edges connecting the data points to the centroids to be $1$ if the distance is less than or equal to $\eta_{d}$, and 00 otherwise, essentially removing the edge.
-This prevents data points from being assigned to centroids that are further than $\eta_{d}$ away, even if it means that they do not get assigned to any cluster.
-Intuitively, this constraint has several benefits.
-First, it provides a form of resistance to outliers, as data points that are far away from any centroid will not (and in some cases, should not) be assigned to any cluster.
-Second, it allows for more fine-grained control over the clusters, as the maximum distance constraint can be used to ensure that the clusters are compact.
-Most importantly, the maximum distance constraint helps to tighten the sensitivity bound of the centroids even further.
-Theorem 7.1.
-If constraints (12) are satisfied, then:
-(13) $\Delta^{\mu}\leq\frac{2\eta_{d}}{\eta_{min}}$
-Proof.
-Since the update phase treats every dimension separately (see Section 5.3, we can consider only one dimension.
-Given a previous centroid $\mu_{t-1}$ (that might be randomly initialized), we need to bound the maximum change in the new centroid $\mu_{t}$, when changing the neighbouring dataset.
-Specifically, we bound the following
-(14) $\Delta^{\mu}=\max\limits_{D\sim D^{\prime}\in\mathcal{D}}||\mu_{t}(D,\mu_{t-1})-\mu_{t}(D^{\prime},\mu_{t-1})||_{1}.$
-In the unbounded model, $D\sim D^{\prime}$ means that the datasets are identical except for one data point, $x^{\prime}$, which is in $D^{\prime}$ but not in $D$.
-If $x^{\prime}$ is not within $\eta_{d}$ of $\mu_{t-1}$, it will not be assigned to the new cluster, and the neighbouring centroids will be the same, completing the proof.
-Therefore, we only need to consider the case where $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$.
-In this case, we have two neighbouring assignments, $O_{t}$ and $O_{t}^{\prime}=O_{t}\cup\{x^{\prime}\}$, and the corresponding centroids will be:
-$\mu_{t}$ $=$ $\frac{\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|}$
-$\mu_{t}^{\prime}$ $=$ $\frac{x^{\prime}+\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|+1}$
-Now, letting $\Sigma=\sum_{x_{l}\in O_{t}}x_{l}$, and $C=|O_{t}|$ to simplify the notation, we bound the difference between the neighbouring centroids as follows:
-(15) $||\mu_{t}-\mu_{t}^{\prime}||_{1}$ $=\left|\left|\frac{\Sigma}{C}-\frac{\Sigma+x^{\prime}}{C+1}\right|\right|_{1}=\left|\left|\frac{\Sigma-Cx_{l}}{C(C+1)}\right|\right|_{1}$
-(16) $\leq\left|\left|\frac{\Sigma-C(\mu_{t}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
-(17) $=\left|\left|\frac{\Sigma-C(\frac{\Sigma}{C}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
-(18) $=\left|\left|\frac{\mp 2\eta_{d}C}{C(C+1)}\right|\right|_{1}\leq\frac{2\eta_{d}}{C}\leq\frac{2\eta_{d}}{\eta_{min}}$
-where (18) follows from the fact that $C\geq\eta_{min}$ and (16) follows from the fact that $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$, and therefore by the triangle inequality:
-$||x_{l}-x^{\prime}||_{2}\leq||x_{l}-\mu_{t-1}||_{2}+||\mu_{t-1}-x^{\prime}||_{2}\leq 2\eta_{d}\quad\forall x_{l}\in O_{t}$
-which implies that $x^{\prime}$ is within $2\eta_{d}$ of any point in $O_{t}$, namely their average $\mu_{t}$.
-Since the $\ell_{1}$ norm over one dimension ($\ell_{\infty}$) is always less than or equal to the $\ell_{2}$ norm, we have:
-$||x^{\prime}-\mu_{t}||_{1}\leq||x^{\prime}-\mu_{t}||_{2}\leq 2\eta_{d}$
-implying that, in this dimension, $x^{\prime}=\mu_{t}\pm 2\eta_{d}$ in the worst case. ∎
-We heuristically set this parameter as follows (for 2 dimensions):
-(19) $\eta_{d}=\sqrt{\frac{B^{2}}{2t}}$
-The intuition behind this choice is that it allows a larger distance ($\eta_{d}\approx 0.7B$) at the beginning of the algorithm (to allow the initial centroids to align with the data) and then gradually reduces the distance, to reduce the noise and fine-tune the clusters.
-In particular, we propose that $\eta_{d}$ should be a function of $k$ and $d$, as the higher number of clusters implies the more compact they are expected to be, and the higher number of dimensions implies that the data points are likely to be more spread out.
-A preliminary evaluation using the S1 and Birch2 datasets shows that this constraint can significantly improve the utility of the algorithm.
-Figure 7 shows this improvement on both the constrained sum-count method CSuLloyd and the constrained centroid method FastLloyd, where we observe a $30\sim 40\%$ reduction in NICV in small $\epsilon$.
-The trade-off with this maximum radius constraint is that the network flow becomes harder to satisfy.
-We find that for other datasets, such as Adult, the min-cost flow solver does not always converge.
-We leave further exploration of this constraint to future work.
-7.2. Serverless Extensions
-In this section, we discuss alternatives to using the service provider.
-A trivial alternative is that each client could locally noise their centroids before sending them to another client (or set of clients) to aggregate.
-The downside to this approach is that the error will compound with the number of clients and decrease utility.
-An alternative is for all or a subset of the clients to take the role of a computation node in an MPC protocol.
-In this case, the centroids would be secret shared and aggregated in MPC, then Laplace noise would need to be added similar to FastLloyd.
-The noise addition can be implemented by each computation node sampling noise locally and adding it to their shares.
-Depending on how many nodes are assumed to be honest, this will necessarily add more noise than FastLloyd to ensure privacy in the presence of one or more nodes colluding.
-Alternatively, the computation nodes can jointly sample the noise using the protocol of Wu et al.
-(Wu et al., 2016).
-The disadvantage of this approach is that it will add a large computational overhead which will make the runtime closer to that of Mohassel et al.
-(Mohassel et al., 2020).
-8. Conclusion
-In this work, we designed FastLloyd, a new private protocol for federated $k$-means.
-We have shown that FastLloyd is secure in the computational model of DP and analyzed its utility.
-Compared to state-of-the-art solutions in the central model of DP, FastLloyd results in higher utility across a wide range of real datasets and scales effectively to larger dimensions and number of clusters.
-FastLloyd also achieves four orders of magnitude faster runtime than the state-of-the-art in secure federated $k$-means across a variety of problem sizes.
-In summary, we provide an efficient, private, and accurate solution to the horizontally federated $k$-means problem.
-Acknowledgements.
-We gratefully acknowledge the support of the National Research Council Canada, the University of Waterloo, the Government of Ontario, and the NSERC Postgraduate Scholarship-Doctoral program for funding this research.
-Appendix A Proof of Quantized DP Noise
+A.3 Proof of Quantized DP Noise
 We argue why quantizing the DP noise is acceptable in Figure 1.
-In essence, because the sensitive data is already quantized, adding quantized noise to it is equivalent to adding the noise first and then performing the quantization, which aligns with the postprocessing lemma in differential privacy.
+In essence, because the sensitive data is already quantized, adding quantized noise to it is equivalent to adding the noise first and then performing the quantization, which aligns with the post-processing lemma in differential privacy.
 This can be intuitively understood by noting that:
@@
 Proof.
-We begin by observing that any value, say $v$, can be split into two parts upon quantization: the integral part, $\bar{v}$, and the fractional part, $\hat{v}$, such that $\bar{v}\times\mathbf{2^{\delta}}$ is the (rounded) integral part of $v\times\mathbf{2^{\delta}}$ (i.e.
-it is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$), and $\hat{v}$ denotes the fractional part that gets lost due to quantization.
+We begin by observing that any value, say $v$, can be split into two parts upon quantization: the integral part, $\bar{v}$, and the fractional part, $\hat{v}$, such that $\bar{v}\times 2^{q}$ is the (rounded) integral part of $v\times 2^{q}$ (i.e.
+it is $\lceil{v\times 2^{q}}\rfloor$), and $\hat{v}$ denotes the fractional part that gets lost due to quantization.
 This notation can similarly be applied to the noise variable, $\eta$.
@@
 We aim to demonstrate the following equivalence:
-$\tilde{v}+\tilde{\eta}=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta\times\mathbf{2^{\delta}}}\rfloor=\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor$
-The right-hand side represents quantization applied as postprocessing to the sum of the sensitive data and the noise, which adheres to the rules of differential privacy.
+$\tilde{v}+\tilde{\eta}=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta\times 2^{q}}\rfloor=\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor$
+The right-hand side represents quantization applied as post-processing to the sum of the sensitive data and the noise, which adheres to the rules of differential privacy.
 After expanding the right-hand side, we obtain:
-$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{\bar{v}\times\mathbf{2^{\delta}}+\bar{\eta}\times\mathbf{2^{\delta}}+\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
-We can take out $\bar{v}\times\mathbf{2^{\delta}}$ and $\bar{\eta{}}\times\mathbf{2^{\delta}}$ from the rounding operator because they are exact integers (indeed $\bar{v}\times\mathbf{2^{\delta}}$ is $\lceil{v\times\mathbf{2^{\delta}}}\rfloor$).
+$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{\bar{v}\times 2^{q}+\bar{\eta}\times 2^{q}+\hat{\eta}\times 2^{q}}\rfloor$
+We can take out $\bar{v}\times 2^{q}$ and $\bar{\eta{}}\times 2^{q}$ from the rounding operator because they are exact integers (indeed $\bar{v}\times 2^{q}$ is $\lceil{v\times 2^{q}}\rfloor$).
 We then obtain:
-$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor+\lceil{\hat{\eta}\times\mathbf{2^{\delta}}}\rfloor$
-However, $\hat{\eta}\times\mathbf{2^{\delta}}$ will be lost due to quantization, hence:
-$\lceil{(\bar{v}+\eta)\times\mathbf{2^{\delta}}}\rfloor=\lceil{v\times\mathbf{2^{\delta}}}\rfloor+\lceil{\eta{}\times\mathbf{2^{\delta}}}\rfloor=\tilde{v}+\tilde{\eta}$
+$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta{}\times 2^{q}}\rfloor+\lceil{\hat{\eta}\times 2^{q}}\rfloor$
+However, $\hat{\eta}\times 2^{q}$ will be lost due to quantization, hence:
+$\lceil{(\bar{v}+\eta)\times 2^{q}}\rfloor=\lceil{v\times 2^{q}}\rfloor+\lceil{\eta{}\times 2^{q}}\rfloor=\tilde{v}+\tilde{\eta}$
 Appendix B Ablation of Parameters
-Figure 9. Ablation over different selections of parameters, $\epsilon=1.0$
-Figure 10. Ablation over different selections of $\alpha$, $\epsilon=0.1$
-In this section, we empirically justify our choice of hyperparameters: $\eta_{min}$, $\eta_{max}$ and the postprocessing method.
+Figure 5: Ablation of $\alpha$ and post-processing strategy on NICV for $\epsilon=0.1$.
+In this section, we empirically justify our choice of hyperparameter $\alpha$ and the post-processing strategy.
 We choose these parameters to minimize the Normalized Intra-Cluster Variance (NICV) metric over a set of synthetic datasets.
-The synthetic datasets were generated using the clusterGeneration package (Qiu and Joe, 2023) varying the number of dimensions ($d$) and clusters ($k$) as $(k,d)\in\{2,4,8,16\}^{2}$, resulting in 16 variations.
-For each variation, we generate 3 different datasets with the degrees of separation $\{0.25,0.5,0.75\}$.
+The synthetic datasets were generated using the clusterGeneration package [58] varying the number of dimensions ($d$) and clusters ($k$) as $(k,d)\in\{2,4,8,16\}^{2}$, resulting in 16 variations.
+For each variation, we generate three different datasets with the degrees of separation $\{0.25,0.5,0.75\}$.
 The average cluster size is fixed to $C_{avg}=\frac{2048}{k}$ and the cluster sizes are randomly sampled in the range $[0.70\cdot C_{avg},1.30\cdot C_{avg}]$, leading to a total dataset size in the range $[1433,2662]$.
-All the ablation experiments were done on two privacy budget settings: $\epsilon=1.0$ and $\epsilon=0.1$.
-The results are shown in Figures 9 and 10, respectively.
-For the min and max cluster experiments, we plot the average of all the NICV values for each $k$ as a separate line and the average of all datasets over all variations as the black dotted line, with the 95% confidence interval shaded around it.
-We observe that generally, a smaller $\alpha_{min}$ and $\alpha_{max}$ lead to better NICV values, especially for datasets with a high number of clusters.
-This is expected because a smaller $\alpha_{min}$ gives a tighter bound on the sensitivity and so improves the utility.
-A smaller $\alpha_{max}$ also imposes a tighter maximum cluster size constraint, which intuitively helps with the convergence of the algorithm.
-However, the tighter the constraints, the harder it is for them to be satisfied.
-As we observe in the plots, most datasets with $\alpha_{min}$ or $\alpha_{max}$ set to lower than $1.25$ do not have a value because the min-cost flow solver cannot find a valid assignment.
-Therefore, we choose $\alpha_{min}=1.25$ and $\alpha_{max}=1.25$ as a balance between utility and feasibility.
-We also observe that the folding method is better in the low privacy setting ($\epsilon=0.1$), and it does not make a significant difference in the high privacy setting ($\epsilon=1.0$).
-Therefore, we choose the folding method as the default postprocessing method.
+All the ablation experiments were done on $\epsilon=0.1$.
+The results are shown in Figure 5, where we plot the average NICV over all datasets for different values of $\alpha$ and post-processing strategies.
+The average is then divided by the NICV of the naive baseline of no post-processing and no radius constraint.
+We observe that the “folding” post-processing strategy consistently outperforms the “truncation" and “none” strategies.
+We also observe that the “Step” strategy outperforms the “Constant” strategy and provides most performance improvement for $\alpha\approx 0.8$.
+We also note that “radius clipping” provides a consistent improvement.
+Based on these results, we choose $\alpha=0.8$ and the “Step” constraint strategy with a post-processing strategy of “folding" after “radius clipping” is applied.
+B.1 Error Analysis of Simple Gaussian Mechanism
+We follow a similar analysis to Section 5.4 to derive the parameters used for GLloyd.
+Since we do not modify the algorithm (only change the noise distribution), we can use Su et al.’s MSE analysis unchanged:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{N^{2}}\left(\mathsf{Var}\left(\gamma^{S}\right)+4\rho^{2}\mathsf{Var}\left(\gamma^{C}\right)\right)$ (28)
+Substituting the variance following our noise mechanism (with the domain-based sensitivity) gives:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{N^{2}}\left(dT(\sigma^{S})^{2}+4\rho^{2}T(\sigma^{C})^{2}\right).$ (29)
+Minimizing this using Lagrange multipliers such that:
+$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (30)
+gives the following ratio of noise multipliers:
+$\sigma^{C}=\sqrt{\frac{\sqrt{d}}{2\rho}}\sigma^{S}$ (31)
+Substituting this back into the MSE equation gives:
+$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}dT\sigma^{2}(2\rho+\sqrt{d})^{2}}{N^{2}}$ (32)
+Setting the per iteration MSE to be less than $0.004$ and rearranging for $T$ gives:
+$T\leq\frac{N^{2}(0.004)}{k^{3}dT\sigma^{2}(2\rho+\sqrt{d})^{2}}$ (33)
+which we also truncate to be in $[2,7]$.
+Appendix C Serverless Extensions
+In this section, we discuss alternatives to using the service provider.
+A trivial alternative is that each client could locally noise their centroids before sending them to another client (or set of clients) to aggregate.
+The downside to this approach is that the error will compound with the number of clients and decrease utility.
+An alternative is for all or a subset of the clients to take the role of a computation node in an MPC protocol.
+In this case, the centroids would be secret shared and aggregated in MPC, then Laplace noise would need to be added similar to FastLloyd.
+The noise addition can be implemented by each computation node sampling noise locally and adding it to their shares.
+Depending on how many nodes are assumed, to be honest, this will necessarily add more noise than FastLloyd to ensure privacy in the presence of one or more nodes colluding.
+Alternatively, the computation nodes can jointly sample the noise using the protocol of Wu et al.
+[68].
+The disadvantage of this approach is that it will add a large computational overhead.
+However, it would still be significantly faster than related work such as Mohassel et al.
+[50], which would incur a similar overhead in addition to their already high computation cost.